### v3

# Enveloping Ellis semigroups as compactifications of transformations groups

Abstract.
The notion of a proper Ellis semigroup compactification is introduced. Ellis’s functional approach shows how to obtain them from totally bounded equiuniformities on a phase space $X$ when the acting group $G$ is with the topology of pointwise convergence and the $G$-space $(G,X,\curvearrowright)$ is $G$-Tychonoff.
The correspondence between proper Ellis semigroup compactifications of a topological group and special totally bounded equiuniformities (called Ellis equiuniformities) on a topological group is established. The Ellis equiuniformity on a topological transformation group $G$ from the maximal equiuniformity on a phase space $G/H$ in the case of its uniformly equicontinuous action is compared with Roelcke uniformity on $G$.
Proper Ellis semigroup compactifications are described for groups ${\rm S}(X)$ (the permutation group of a discrete space $X$) and $\mbox{{\rm Aut}}(X)$ (automorphism group of an ultrahomogeneous chain $X$) in the permutation topology. It is shown that this approach can be applied to the unitary group of a Hilbert space.
Keywords: semigroup, topological transformation group, enveloping Ellis semigroup, equiuniformity, compactification, ultratransitive action, chain, Hilbert space.
AMS classification: Primary 57S05, 20E22 Secondary 22F05, 22F50, 54E05, 54D35, 54H15, 47B02
The study of the first author was carried out with the help of the Center of Integration in Science, Ministry of Aliyah and Integration, Israel, 8123461 and is supported by ISF grant 3187/24.
1. Introduction and preliminary remarks
Compactifications of topological (semi)groups are a vital tool of investigations in Topological dynamics [47], [19] and Topological Algebra [21]. In the paper by a proper compactification of a topological space $X$, we understand a compactum in which $X$ is a dense subspace. Due to A. Weil a topological group $G$ is a dense subgroup of a compact topological group $K$ ($K$ is a proper compactification of $G$) iff $G$ is totally bounded (in the right uniformity). There are groups without (proper) semitopological semigroup compactifications [32]. Every topological group has a proper right topological semigroup compactification, the greatest ambit [7]. A self-contained treatment of the theory of compact right topological semigroups and, in particular, of semigroup compactifications (nonproper) is in [6]. A study of the proper group’s compactifications can be found in [46].
The functional approach in the study of transformation groups (groups acting effectively on phase spaces) allows us to view elements of $G$ as self maps of the phase space $X$. Each $g\in G$ is an element of $X^{X}$ and the composition of maps makes the image of $G$ a semigroup. Moreover, the multiplication on $G$ and composition on $X^{X}$ makes this representation an isomorphism of semigroups. This approach came from topological dynamics due to Ellis’s construction [13]. R. Ellis examined the subgroup $G$ of group of homeomorphisms $\mbox{{\rm Hom}}(X)$ of compactum $X$ as a subsemigroup of the Cartesian product $X^{X}$ (a right topological semigroup) and named the closure of $G$ in $X^{X}$ the enveloping (Ellis) semigroup [13]. It is a compact right topological semigroup and is a nonproper semigroup compactification of $G$ ($G$ is not its topological subspace).
However, if $G$ is in the topology of pointwise convergence (for the action $G\curvearrowright X$), then Ellis’s representation becomes a topological isomorphism. For the action of a topological group on itself by multiplication on the left the topology of a group is equal to the topology of pointwise convergence (Lemma 1.3). If the topology of pointwise convergence on a group is an admissible group topology for the action $G\curvearrowright X$ and $(G,X,\curvearrowright)$ is a $G$-Tychonoff space, then the topology of pointwise convergence on a group is an admissible group topology for the action of $G$ of any $G$-compactification of $X$ (Lemma 1.4). These observations allow us to define in § 2.1 a concept of a proper Ellis semigroup compactification (Definition 2.1) and the corresponding Ellis equiuniformity (Definition 2.3). The proper Ellis semigroup compactification is a right topological semigroup and is a $G$-compactification of a group $G$ (multiplication agrees with the extended action). We demonstrate how to obtain all Ellis equiuniformities from totally bounded equiuniformities on $G$ ($G$ is examined as a $G$-Tychonoff space $(G,G,\cdot)$) in § 2.4 (Theorem 2.22).
In § 3 we examine uniformly equicontinuous actions (in this case the topology of pointwise convergence is an admissible group topology on an acting group, and actions by isometries and actions on coset spaces with respect to neutral subgroups (by multiplication on the left) are among them). For actions on coset spaces with respect to a neutral subgroup, the corresponding Ellis equiuniformity is less than or equal to the least Ellis equiuniformity on a group that is greater than or equal to the precompact reflection of the Roelcke uniformity (Theorem 3.2).
Examples illustrate what proper Ellis semigroup compactifications can be obtained from the totally bounded maximal equiuniformities on phase spaces.
In § 3.1.1 we examine an example of an ultratransitive action of a group $G$ on a discrete space $X$. It is shown that the Roelcke compactification of $G$ is equal to the proper Ellis semigroup compactification obtained from the maximal equiuniformity on $X$ and is a semitopological semigroup. The maximal equiuniformity in this case is the least totally bounded uniformity on $X$ (and corresponds to the Alexandroff compactification of $X$). The Roelcke compactification of $G$ is described (Theorem 3.4 and Corollary 3.8). Houghton’s groups serve as an example. In [18, § 12, Theorem 12.2] the description of the Roelcke compactification of the permutation group of natural numbers is given. It is also shown that it is a semitopological group homeomorphic to the Cantor set.
In § 3.1.2 we examine an ultratransitive action of a subgroup $G$ of automorphism group $\mbox{{\rm Aut}}(X)$ on a chain $X$ in discrete topology. We establish that the Roelcke compactification of $G$ is a proper Ellis semigroup compactification iff a chain is continuous (Corollaries 3.13 and 3.17). In the general case, we derive the least Ellis equiuniformity which is greater than or equal to the Roelcke uniformity, from the Roelcke uniformity using Ellis construction (Theorem 3.11 and Corollary 3.17). The proper Ellis semigroup compactification and the Roelcke compactification of $G$ are described in Theorem 3.11. The maximal equiuniformity on $X$ from which the Ellis equiuniformity is obtained corresponds to the maximal $G$-compactification of $X$ which is also described in § 3.1.2. It is the minimal linearly ordered compactification of a GO-space $X$. Thompson group $F$ in the permutation topology is examined in Example 3.18. $F$ is Roelcke precompact but its Roelcke compactification is not a proper Ellis semigroup compactification.
In § 3.2 it is shown that the Roelcke uniformity on the unitary group of a Hilbert space coincides with the uniformity from the family of the stabilizers of points of the unit sphere and is an Ellis equiuniformity (Theorem 3.23). The result provides an alternative proof that the Roelcke compactification of the unitary group is a semitopological group which coincides with the space of all linear operators of norm 1$\leq$ 1≤ 1 in the weak operator topology [45].
The method for working with Ellis equiuniformities is detailed in § 2.2 and 2.3. In Theorem 2.7 we present both the topological and uniform variants of Ellis’s construction and establish the extremal property of Ellis uniformity. Proposition 2.14 outlines a condition for the coincidence of the Ellis equiuniformity corresponding to the extended action and equiuniformity corresponding to the original one. We demonstrate that the map of the poset of equiuniformities on the phase space to the poset of Ellis equiuniformities preserves order. In Proposition 2.16 we establish a condition when the Ellis equiuniformity is equal to the uniformity on a group from a family of small subgroups. Lastly, in Proposition 2.21 we establish the connection between Ellis equiuniformities for the actions on different phase spaces.
The paper is a continuation of investigations initiated in [42].
We adopt terminology and notation from [14] and [37]. The family of finite subsets of a set (or space) $X$ is denoted $\Sigma_{X}$ and directed by inclusion. The set of non-negative reals is represented by $\mathbb{R}_{+}$.
[…]
Uniform space is denoted as $(X,\mathcal{U})$, $u$ is a cover and ${\rm U}$ is a corresponding entourage of $\mathcal{U}$. If the cover $v$ is a refinement of $u$, we use the notation $v\succ u$. Uniformity on a topological space is compatible with its topology. Additionally, $(\tilde{X},\tilde{\mathcal{U}})$ denotes the completion of $(X,\mathcal{U})$.
1.1. Semigroups
A semigroup $(S,\bullet)$ is a set $S$ with an associative internal binary operation (multiplication $\bullet$). A group $G$ with multiplication $\cdot$ is a semigroup with multiplication $\bullet$=$\cdot$ = ⋅. A map $s:(S,\bullet)\to(S^{\prime},\bullet^{\prime})$ of semigroups is a homomorphism if $s(x\bullet y)=s(x)\bullet^{\prime}s(y)$ for all $x,y\in S$. If $s:(G,\cdot)\to(S,\bullet)$ is a homomorphism, then $s(G)$ is a group and a subsemigroup of $S$.
A semitopological (right topological) semigroup is a semigroup $(S,\bullet)$ on a topological space $S$ with separately continuous multiplication $\bullet$ (multiplication $\bullet$ continuous on the right).
All the necessary (and additional) information can be found in [6] and [3].
1.2. Spaces of maps and semigroup structure on $X^{X}$
1. Since the Cartesian product $X^{X}$ may be regarded as the set of maps of $X$ into $X$ ($f=(f(x))_{x\in X}\in X^{X}$), a semigroup structure (multiplication is composition $\circ$ of maps) may be introduced on $X^{X}$.
If $X$ is a topological space, then the multiplication on $X^{X}$ in the Tychonoff topology (the subbase is formed by the sets $[x,O]=\{f\in X^{X}\ |\ f(x)\in O\}$, $x\in X$, $O$ open subset of $X$) is continuous on the right ($f\to f\circ g$; if $f\in[g(x),O]$, then $f\circ g\in[x,O]$) and continuous on the left for continuous maps ($f\to g\circ f$; $g\in C(X)$, if $f\in[x,g^{-1}O]$, then $g\circ f\in[x,O]$). Thus, $(X^{X},\circ)$ is a right topological semigroup.
If $(X,\mathcal{U}_{X})$ is a uniform space, then the uniformity $\mathcal{U}_{X}^{p}$ on $X^{X}$ is the Cartesian product of uniformities $\mathcal{U}_{X}$ on factors (see [14, Ch. 8, § 8.2]). The multiplication on $(X^{X},\mathcal{U}_{X}^{p})$ is uniformly continuous on the right ($f\to f\circ g$; if $y=g(x)$ and $(f_{1}(y),f_{2}(y))\in U\in\mathcal{U}_{X}$, then $((f_{1}\circ g)(x),(f_{2}\circ g)(x))\in U$) and uniformly continuous on the left for uniformly continuous maps ($f\to g\circ f$, $g$ is uniformly continuous; for ${\rm U}\in\mathcal{U}_{X}$ let ${\rm V}\in\mathcal{U}_{X}$ be such that if $(x,y)\in{\rm V}$, then $(g(x),g(y))\in{\rm U}$; thus if $(f_{1}(x),f_{2}(x))\in{\rm V}$, then $((g\circ f_{1})(x),(g\circ f_{2})(x))\in{\rm U}$).
The topology on $X^{X}$ induced by $\mathcal{U}_{X}^{p}$ is the Tychonoff topology of the product where every factor has the Tychonoff topology induced by $\mathcal{U}_{X}$. Following [9, Ch.10 , § 1] the uniformity $\mathcal{U}_{X}^{p}$ on $X^{X}$ (as the set of maps) is called the uniformity of pointwise (or simple) convergence (abbreviation u.p.c) and the induced topology is the topology of pointwise convergence $\tau_{p}$ (abbreviation t.p.c.) [14, Ch. 2, § 2.6].
2. If $(X,\mathcal{U}_{X})$ is a uniform space, then, following [9, Ch.10 , § 1] (see, also [14, Ch. 8, § 8.2]), the uniformity of uniform convergence $\mathcal{U}_{X}^{u}$ on $X^{X}$ (abbreviation u.u.c.) is defined. The topology $\tau_{u}$ on $X^{X}$ induced by $\mathcal{U}_{X}^{u}$ is the topology of uniform convergence (abbreviation t.u.c.). Evidently, ${\mathcal{U}}_{X}^{p}\subset\mathcal{U}_{X}^{u}$ and $\tau_{u}\geq\tau_{p}$.
1.3. $G$-spaces, admissible group topologies on a transformation group and the topology of pointwise convergence
1. An action $\alpha:G\times X\to X$ (designation $\alpha(g,x)=gx$) of a group $G$ on a set $X$ (phase space) is called effective if the kernel of the action $\{g\in G\ |\ gx=x,\ \forall\ x\in X\}$ is the unit of $G$. If $G$ effectively acts on $X$, then $G\subset{\rm S}(X)$, where ${\rm S}(X)$ is the permutation group of $X$.
The subgroup $\mbox{{\rm St}}_{\sigma}=\{g\in G\ |\ gx=x,\ x\in\sigma\in\Sigma_{X}\}=\bigcap\limits_{x\in\sigma}\mbox{{\rm St}}_{x}$ of $G$ is a stabilizer (or stabilizer of points from $\sigma$).
If $T\subset G$, $Y\subset X$, then $TY=\alpha(T,Y)=\bigcup\{ty\ |\ t\in T,\ y\in Y\}$. For $x\in X$ $\alpha_{x}:G\to X$, $\alpha_{x}(g)=\alpha(g,x)$, is an orbit map.
2. Under a continuous action $\alpha:G\times X\to X$ of a topological group $G$ on a space $X$, the triple $(G,X,\alpha)$ is called a $G$-space. A continuous action $\alpha:G\times X\to X$ will be denoted $G\curvearrowright X$ if we don’t need the special designation for the map. $G$-spaces $(G,X,\alpha)$ and $(G,X^{\prime},\alpha^{\prime})$ are equivalent if there exists a homeomorphism $\varphi:X\to X^{\prime}$ such that $\varphi(\alpha(g,x))=\alpha^{\prime}(g,\varphi(x))$. Equivalent $G$-spaces are identified.
A uniformity $\mathcal{U}_{X}$ on $X$ is called an equiuniformity [29] if the action $G\curvearrowright X$ is saturated (any homeomorphism from $G$ is uniformly continuous) and is bounded (for any $u\in\mathcal{U}$ there are $O\in N_{G}(e)$ and $v\in\mathcal{U}$ such that the cover $Ov=\{OV\ |\ V\in v\}\succ u$). In this case $(G,X,\alpha)$ a $G$-Tychonoff space, the action is extended to the continuous action $\tilde{\alpha}:G\times\tilde{X}\to\tilde{X}$ on the completion $(\tilde{X},\tilde{\mathcal{U}}_{X})$ of $(X,\mathcal{U}_{X})$. The extension $\tilde{\mathcal{U}}_{X}$ of $\mathcal{U}_{X}$ to $\tilde{X}$ is an equiuniformity on $\tilde{X}$.
The embedding $\jmath:X\to\tilde{X}$ is a $G$-map or an equivariant map of $G$-spaces, i.e. that the following diagram
$\begin{array}[]{ccc}\quad G\times X&\stackrel{{\scriptstyle{\mbox{{\rm id}}\times\jmath}}}{{\longrightarrow}}&G\times\tilde{X}\\ \alpha\downarrow&&\quad\downarrow\tilde{\alpha}\\ \quad X&\stackrel{{\scriptstyle\jmath}}{{\longrightarrow}}&\tilde{X},\end{array}$
is commutative, the pair $(\tilde{X},\jmath)$ is a $G$-extension of $X$ and $\jmath(X)$ is a dense invariant subset of $\tilde{X}$ [31]. $G$-spaces as extensions will also be denoted as $(G,(\tilde{X},\jmath),\tilde{\alpha})$.
If $\mathcal{U}_{X}$ is a totally bounded equiuniformity on $X$, then $(bX=\tilde{X},\jmath_{b})$ is a $G$-compactification or an equivariant compactification of $X$. If $(G,X,\curvearrowright)$ is a $G$-Tychonoff space, then the maximal (totally bounded) equiuniformity on $X$ exists and the maximal $G$-compactification $(\beta_{G}X,\jmath_{\beta})$ of $X$ corresponds to the maximal totally bounded equiuniformity on $X$. The maximal totally bounded equiuniformity $\mathcal{U}_{X}^{bm}$ on $X$ is the precompact reflection [23] of the maximal equiuniformity $\mathcal{U}_{X}^{m}$ on $X$ and the compactification $(bX,\tilde{\mathcal{U}}_{X}^{bm})$ (completion of $(X,{\mathcal{U}}_{X}^{bm})$) is the Samuel compactification of $X$ with respect to the equiuniformity $\mathcal{U}_{X}^{m}$ (see [14, Ch. 8, Problem 8.5.7]).
Remark 1.1.
If for an action $G\curvearrowright X$ (not continuous) of a topological group $G$ on a space $X$ the uniformity $\mathcal{U}_{X}$ on $X$ is saturated and bounded, then the the action $G\curvearrowright X$ is continuous.
The maximal equiuniformity on a coset space $G/H$ of a topological group $G$ (action by multiplication on the left) is described in [11] or [25]. There such actions are called open and the covers of the base of the maximal equiuniformity are of the form
$\{Ox\ |\ x\in G/H\},\ O\in N_{G}(e).$
We denote by $\mathbb{BU}(X)$ the set of all totally bounded equiuniformities on $X$. All the necessary (and additional) information can be found in [31].
3. If $X$ is a topological space, $\mbox{{\rm Hom}}(X)$ is the group of its homeomorphisms, then a group $G$ effectively acts on $X$ if $G\subset\mbox{{\rm Hom}}(X)$. If $X$ is a discrete space, then $\mbox{{\rm Hom}}(X)={\rm S}(X)$.
A topology in which a group $G$ is a topological group and its effective action $G\curvearrowright X$ is continuous is called an admissible group topology on $G$ [1].
For an effective action of $G$ on a compactum $X$ the compact-open topology $\tau_{co}$ is the smallest admissible group topology on $G$ [1]. Considering elements of $G$ as continuous maps of $X$ into $X$ ($G\subset X^{X}$) the compact-open topology $\tau_{co}$ coincides with the t.u.c. $\tau_{u}$ [14, Corollary 8.2.7].
For an effective action of $G$ on a discrete space $X$, the permutation topology $\tau_{\partial}$ (the subbase is formed by the sets $[x,y]=\{g\in G\ |\ gx=y\}$, $x,y\in X$) is the smallest admissible group topology. The group $G$ in the permutation topology is non-Archimedean (a unit nbd base is formed by clopen subgroups (stabilizers)) and the $G$-space $((G,\tau_{\partial}),X,\curvearrowright)$ is $G$-Tychonoff.
If the topology of pointwise convergence $\tau_{p}$ (the subbase is formed by sets of the form $[x,O]=\{g\in G\ |\ gx\in O\}$, $x\in X$, $O$ is open in $X$) is an admissible group topology on a group $G$ of homeomorphisms of the space $X$ acting effectively, then it is the smallest admissible group topology [26, Lemma 3.1], $\tau_{\partial}\geq\tau_{p}$, $\tau_{\partial}$ is an admissible group topology and $\tau_{\partial}=\tau_{p}$, if $X$ is a discrete space. Considering elements of $G$ as continuous maps of $X$ into $X$ ($G\subset X^{X}$) the topology of pointwise convergence $\tau_{p}$ coincides with the t.p.c. from § 1.2 [14, Proposition 2.6.3].
If the topology of pointwise convergence $\tau_{p}$ is an admissible group topology on $G$ acting effectively on a compactum $X$, then $\tau_{p}=\tau_{co}$.
Remark 1.2.
Example from [30] shows that the t.p.c. $\tau_{p}$ may be an admissible group topology on a group $G$ of homeomorphisms of $X$ acting effectively, but the $G$-space $(G=(G,\tau_{p}),X,\curvearrowright)$ may not be $G$-Tychonoff.
In the paper, effective actions are examined.
4. Let $G$ be a topological group. The $G$-space $(G,G,\cdot)$ with the action $G\curvearrowright G$ by multiplication $\cdot$ on the left is $G$-Tychonoff.
Lemma 1.3.
Let $(G,\tau)$ be a topological group. For the $G$-space $(G,G,\cdot)$ the t.p.c. $\tau_{p}$ is an admissible group topology on $G$ and $\tau=\tau_{p}$.
Proof.
[…]
Lemma 1.4.
Let $(G=(G,\tau_{p}),X,\alpha)$ be $G$-Tychonoff and $\mathcal{U}_{X}$ is an equiuniformity on $X$. Then the t.p.c. $\tilde{\tau}_{p}$ on $G$ for the extended action $\tilde{\alpha}:G\times\tilde{X}\to\tilde{X}$ ($(\tilde{X},\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$) is the smallest admissible group topology on $G$ and $\tilde{\tau}_{p}=\tau_{p}$.
Proof.
Let $\mathcal{V}_{X}$ be the precompact reflection of $\mathcal{U}_{X}$, $(bX,\tilde{\mathcal{V}}_{X})$ is the Samuel compactification of $(X,\mathcal{U}_{X})$ and $(bX,\jmath_{\mathcal{V}_{X}})$ is the $G$-compactification of $X$. From the correctly defined commutative diagram
$\begin{array}[]{ccl}X&\stackrel{{\scriptstyle\jmath_{\mathcal{V}_{X}}}}{{\hookrightarrow}}&\quad bX\\ \quad\jmath_{\mathcal{U}_{X}}\searrow&&\nearrow\jmath\\ &\tilde{X}&\end{array}$
where all maps are equivariant embeddings and $\jmath$ is the extension of a uniformly continuous map $\jmath_{\mathcal{V}_{X}}$ (see [14, Ch. 8, Problem 8.5.7]), it follows that $\jmath_{\mathcal{V}_{X}}(X)(=X)$ and $\jmath(\tilde{X})(=\tilde{X})$ are invariant subsets of $bX$.
The compact-open topology $\tau_{co}$ on $G$ induced by the extended action $G\curvearrowright bX$ is the smallest admissible group topology on $G$ and $\tau_{p}$ is an admissible group topology on $G$. Hence, $\tau_{p}\geq\tau_{co}$. Evidently, $\tau_{co}\geq\tau^{s}_{p}$, where $\tau^{s}_{p}$ is the t.p.c. on $G$ for the action $G\curvearrowright bX$ (not necessarily an admissible group topology).
Since $X$ and $\tilde{X}$ are invariant subsets of $bX$, for the t.p.c. $\tilde{\tau}_{p}$ on $G$ for the action $G\curvearrowright\tilde{X}$ (not necessarily an admissible group topology) one has $\tau^{s}_{p}\geq\tilde{\tau}_{p}\geq\tau_{p}$. Therefore, $\tau_{co}=\tilde{\tau}_{p}=\tau_{p}$ and $\tilde{\tau}_{p}$ is the smallest admissible group topology for the extended action $G\curvearrowright\tilde{X}$. ∎
1.4. Uniformities on a topological group
On a topological group $G$ four uniformities are well-known. The right uniformity $R$ (the base is formed by the covers $\{Og=\bigcup\{h\cdot g\ |\ h\in O\}\ |\ g\in G\}$, $O\in N_{G}(e)$), the left uniformity $L$, the two sided uniformity $R\vee L$ (the least upper bound of the right and left uniformities) and the Roelcke uniformity $L\wedge R$ (the greatest lower bound of the right and left uniformities) (the base is formed by the covers $\{OgO=\bigcup\{h\cdot g\cdot h^{\prime}\ |\ h,h^{\prime}\in O\}\ |\ g\in G\}$, $O\in N_{G}(e)$). A group $G$ is Roelcke precompact if the Roelcke uniformity is totally bounded.
For every equiuniformity $\mathcal{U}$ on $G$ of a $G$-space $(G,G,\cdot)$ one has $\mathcal{U}\subset R$, $L\wedge R$ is an equiuniformity. All the necessary information about these uniformities can be found in [37].
Uniformity on $G$ from a family of small subgroups is introduced in [26, § 4]. Let $(G=(G,\tau_{p}),X,\curvearrowright)$ be a $G$-space. The family of stabilizers
$\mbox{{\rm St}}_{\sigma}=\bigcap\{\mbox{{\rm St}}_{x}\ |\ x\in\sigma\},\ \sigma\in\Sigma_{X},$
is a directed family of small (closed) subgroups of $G$ ($\mbox{{\rm St}}_{\sigma}\leq\mbox{{\rm St}}_{\sigma^{\prime}}\Longleftrightarrow\sigma\subset\sigma^{\prime}$) which defines the uniformity $R_{\Sigma_{X}}$ on $G$. The base of $R_{\Sigma_{X}}$ is formed by the covers
[…]
Uniformity $R_{\Sigma_{X}}$ is an equiuniformity (for the action $G\curvearrowright G$) and has a spectral representation. It is initial [37, Proposition-definition 0.16] with respect to the quotient maps $q_{\sigma}:G\to(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})$ (cosets $G/\mbox{{\rm St}}_{\sigma}$ with the maximal equiuniformities $\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}}$ (the quotient uniformities of the right uniformity $R$ on $G$[37, Theorem 5.21] and [11]), $\sigma\in\Sigma_{X}$.
The completion of $G$ with respect to $R_{\Sigma_{X}}$ is uniformly equivalent to the limit of the inverse spectrum $\{\widetilde{G/\mbox{{\rm St}}_{\sigma}},p_{\sigma^{\prime}\sigma},\Sigma_{X}\}$, where $\widetilde{G/\mbox{{\rm St}}_{\sigma}}$ is the completion of $G/\mbox{{\rm St}}_{\sigma}$ with respect to $\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}}$, $p_{\sigma^{\prime}\sigma}:\widetilde{G/\mbox{{\rm St}}_{\sigma^{\prime}}}\to\widetilde{G/\mbox{{\rm St}}_{\sigma}}$ is the extension of the quotient map $G/\mbox{{\rm St}}_{\sigma^{\prime}}\to G/\mbox{{\rm St}}_{\sigma}$, $\sigma\subset\sigma^{\prime}$ [26, Theorem 1.4].
Remark 1.5.
[26, Corollary 4.5] $L\wedge R\subset R_{\Sigma_{X}}\subset R$, hence, if the uniformity $R_{\Sigma_{X}}$ is totally bounded, then $G$ is Roelcke precompact.
For a group $G$ (acting on a discrete space $X$) in the permutation topology the stabilizers $\mbox{{\rm St}}_{\sigma}$, $\sigma\in\Sigma_{X}$, are clopen subgroups (nbds of the unit) and $L\wedge R=R_{\Sigma_{X}}$.
2. Ellis uniformity on a group from an equiuniformity on a phase space
2.1. A proper enveloping Ellis semigroup compactification of a topological group
Definition 2.1.
[…]
$\mathfrak{E}(G)$ is the set of proper enveloping Ellis semigroup compactifications of $G$.
Remark 2.2.
( a) If a topological group $G$ is a subsemigroup of a compact semitopological semigroup $S$, then the restriction of the multiplication to $G\times S$ is necessarily continuous [28, Proposition 6.2], but there are topological groups which have no proper semitopological semigroup compactifications [32].
(b) If in the Definition 2.1 the requirement on $s$ is relaxed to be a continuous homomorphism, then one gets the notion of an enveloping semigroup of $G$ in [47, § 3.6].
(b’) If in the Definition 2.1 $G$ is a semigroup and a topological space, the requirement on $s$ is relaxed to be a continuous homomorphism and the condition of continuity of the restriction of the multiplication to $s(G)\times S$ is relaxed to its continuity on the left, then one gets the notion of a semigroup compactification in [21].
(c) $\mathfrak{E}(G)$ with the order: $(S_{1},s_{1})\leq(S_{2},s_{2})$ if there is a continuous homomorphism $\psi:S_{2}\to S_{1}$ such that $\psi\circ s_{2}=s_{1}$, is a poset. Below it will be shown that the poset $\mathfrak{E}(G)$ is nonempty.
Let $(S,s)\in\mathfrak{E}(G)$. The restriction of the multiplication to $s(G)\times S$ defines a continuous action $\alpha_{S}:G\times S\to S$ ($\alpha_{S}(g,x)=s(g)\bullet x$) and the diagram
$\begin{array}[]{ccccc}\quad G\times G&\stackrel{{\scriptstyle\mbox{{\rm id}}\times s}}{{\hookrightarrow}}&G\times S&\stackrel{{\scriptstyle s\times\mbox{{\rm id}}}}{{\hookrightarrow}}&S\times S\\ \cdot\downarrow&&\quad\downarrow\alpha_{S}&&\quad\downarrow\bullet\\ \quad G&\stackrel{{\scriptstyle s}}{{\hookrightarrow}}&S&\stackrel{{\scriptstyle\mbox{{\rm id}}}}{{\longrightarrow}}&S\end{array}$
is commutative.
The unique uniformity $\tilde{\mathcal{E}}$ on $S$ induces a totally bounded equiuniformity $\mathcal{E}$ on $G$ (the completion of $(G,\mathcal{E})$ is $(S,\tilde{\mathcal{E}})$). Denote by $\mathbb{BU}(G)$ the set of totally bounded equiuniformities on $G$ as a phase space of $(G,G,\cdot)$.
Definition 2.3.
A totally bounded equiuniformity $\mathcal{E}$ on $G$ is an Ellis equiuniformity if $\tilde{G}$ is a proper enveloping Ellis semigroup compactification of $G$ ($(\tilde{G},\tilde{\mathcal{E}})$ is the completion of $(G,\mathcal{E})$).
$\mathbb{E}(G)$ is a poset of Ellis equiuniformities on $G$.
Remark 2.4.
[…]
Proposition 2.5.
The posets $\mathfrak{E}(G)$ and $\mathbb{E}(G)$ are isomorphic.
Proof.
The restriction of isomorphism of posets of compactifications of $X$ and totally bounded uniformities on $X$ [14, Ch.8, § 8.4] to the proper enveloping Ellis semigroup compactifications of $G$ and the Ellis equiuniformities on $G$ will work if one shows that a continuous map $\psi:S_{2}\to S_{1}$ of the proper enveloping Ellis semigroup compactifications $(S_{1},s_{1})\leq(S_{2},s_{2})$ of $G$ such that $\psi\circ s_{2}=s_{1}$ is a homomorphism.
We must check fulfillment of the equality
[…]
From commutativity of the diagram
$\begin{array}[]{ccl}S_{2}&\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}&\quad S_{1}\\ \quad s_{2}\nwarrow&&\nearrow s_{1}\\ &G&\end{array}$
where $s_{2},s_{1}$ are topological isomorphisms, for $x,y\in G$ one has
$\psi(s_{2}(x)\bullet s_{2}(y))=\psi(s_{2}(x\cdot y))=s_{1}(x\cdot y)=s_{1}(x)\bullet s_{1}(y)=\psi(s_{2}(x))\bullet\psi(s_{2}(y))$
and (e1) holds for $(x,y)\in s_{2}(G)\times s_{2}(G)$.
The restrictions of the maps
$\Psi:S_{2}\times S_{2}\to S_{1},\ \Psi(x,y)=\psi(x\bullet y),\ \mbox{and}$
$\Psi^{\prime}:S_{2}\times S_{2}\to S_{1},\ \Psi^{\prime}(x,y)=\psi(x)\bullet\psi(y),$
on the subset $s_{2}(G)\times S_{2}$ are continuous. Hence, (e1) holds for $(x,y)\in s_{2}(G)\times S_{2}$ (since continuous maps coincide on a dense subset $s_{2}(G)\times s_{2}(G)$).
Since $S_{2}$ is a right topological semigroup, for any fixed $y\in S_{2}$ the restrictions of the maps $\Psi$ and $\Psi^{\prime}$ to $S_{2}\times\{y\}$ are continuous and coincide on a dense subset $s_{2}(G)\times\{y\}$. Hence, (e1) holds on $S_{2}\times\{y\}$, $y\in S_{2}$. Thus (e1) holds on $S_{2}\times S_{2}$. ∎
2.2. Diagonal action, agreement of algebraic structures with topology
(A) For a family of sets $X_{t}$ with actions $\alpha_{t}$ of a group $G$, $t\in\mathcal{T}$, the diagonal action $\alpha_{\Delta\mathcal{T}}$ of $G$ on the product $\prod\limits_{t\in\mathcal{T}}X_{t}$
[…]
where $f=(f(t))_{t\in\mathcal{T}}\in\prod\limits_{t\in\mathcal{T}}X_{t}$ (note, that $f(t)=\mbox{{\rm pr}}_{t}(f)$) is correctly defined (see, for example, [47, § 2.7 Constructions, 2]).
Proposition 2.6.
(1) For an action $\alpha:G\times X\to X$ the map
[…]
is an isomorphism of $G$ onto the group $\imath(G)$ which is a subsemigroup of $X^{X}$.
(2) $\alpha_{\Delta X}(g,f)=\imath(g)\circ f$.
(3) The following diagram is commutative ($\imath$ is a $G$-map (not continuous))
$\begin{array}[]{ccc}\quad G\times G&\stackrel{{\scriptstyle{\mbox{{\rm id}}\times\imath}}}{{\hookrightarrow}}&G\times X^{X}\\ \cdot\downarrow&&\quad\downarrow\alpha_{\Delta X}\\ \quad G&\stackrel{{\scriptstyle\imath}}{{\hookrightarrow}}&X^{X}.\end{array}$
Proof.
(1) Since the action is effective $\imath$ is injective. $\forall\ x\in X$
$(\imath(g)\circ\imath(h))(x)=\imath(g)(\imath(h)(x))\stackrel{{\scriptstyle(\imath)}}{{=}}\imath(g)(\alpha(h,x)))\stackrel{{\scriptstyle(\imath)}}{{=}}\alpha(g,\alpha(h,x)))=\alpha(g\cdot h,x)\stackrel{{\scriptstyle(\imath)}}{{=}}\imath(g\cdot h)(x).$
Thus, $\imath(g\cdot h)=\imath(g)\circ\imath(h)$ and on $\imath(G)$ the stucture of a group (with multiplication $\circ$) is correctly defined which is a subsemigroup of $X^{X}$ (see § 1.1 and 1.2).
(2) Let $f=(f(x))_{x\in X}$, $g\in G$. $\forall\ x\in X$
$\alpha_{\Delta X}(g,f)(x)\stackrel{{\scriptstyle(\Delta)}}{{=}}\alpha(g,f(x))\stackrel{{\scriptstyle(\imath)}}{{=}}\imath(g)(f(x))=(\imath(g)\circ f)(x).$
Thus, $\alpha_{\Delta X}(g,f)=\imath(g)\circ f$.
(3) $\forall\ g,h\in G$
$\imath(g\cdot h)\stackrel{{\scriptstyle(1)}}{{=}}\imath(g)\circ\imath(h)\stackrel{{\scriptstyle(2)}}{{=}}\alpha_{\Delta X}(g,\imath(h))=\alpha_{\Delta X}((\mbox{{\rm id}}\times\imath)(g,h)).$
It is easy to check that if $(G,X_{t},\curvearrowright)$ is a $G$-space ($G$-Tychonoff and $\mathcal{U}_{t}$ is the correspondent equiuniformity on $X_{t}$), $t\in\mathcal{T}$, then $(G,\prod\limits_{t\in\mathcal{T}}X_{t},\alpha_{\Delta\mathcal{T}})$ is a $G$-space ($G$-Tychonoff and the Cartesian product $\mathcal{U}^{p}$ of equiuniformities $\mathcal{U}_{t}$, $t\in\mathcal{T}$, is an equiuniformity on $\prod\limits_{t\in\mathcal{T}}X_{t}$).
Theorem 2.7.
I. Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-space. Then
(1) $(X^{X},\circ)$ is a right topological semigroup (besides, multiplication on the left is continuous for $f\in C(X)$),
(2) $\imath:G\to X^{X}$ is a topological isomorphism of $G$ onto the topological group $\imath(G)$ which is a subsemigroup of $X^{X}$ and the restriction of multiplication $\circ$ to $\imath(G)\times X^{X}$ is continuous (besides, the multiplication on the left is continuous for $f\in\imath(G)$, $f\in C(X)$),
(3) $\imath:G\to X^{X}$ is a continuous $G$-map,
(4) the closure $\mbox{{\rm cl}}~\imath(G)$ is an invariant subset of $X^{X}$ (and, hence, $(G,\mbox{{\rm cl}}~\imath(G),\alpha_{\Delta X}|_{G\times\mbox{{\rm cl}}~\imath(G)})$ is a $G$-space) and a subsemigroup of $X^{X}$. The restriction of multiplication $\circ$ to $\imath(G)\times\mbox{{\rm cl}}~\imath(G)$ is continuous (besides, the multiplication on the left is continuous for $f\in\mbox{{\rm cl}}~\imath(G)$, $f\in C(X)$).
II. Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}$ is an equiuniformity on $X$ and $(\tilde{X},\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$. Then
(a) $\mathcal{U}_{X}^{p}|_{\imath(G)}$ is an equiuniformity on $G$, the completion of $(G,\mathcal{U}_{X}^{p}|_{\imath(G)})$ is the closure of $\imath(G)$ in $\tilde{X}^{X}$,
(b) for the $G$-Tychonoff space $(G=(G,\tau_{p}),\tilde{X},\tilde{\alpha})$ $\mathcal{U}_{\tilde{X}}^{p}|_{\tilde{\imath}(G)}$ is an equiuniformity on $G$, the completion of $(G,\mathcal{U}_{\tilde{X}}^{p}|_{\tilde{\imath}(G)})$ is the closure of $\tilde{\imath}(G)$ in $\tilde{X}^{\tilde{X}}$,
(c) $\mathcal{U}_{X}^{p}|_{\imath(G)}\subset\mathcal{U}_{\tilde{X}}^{p}|_{\tilde{\imath}(G)}$.
III. Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}\in\mathbb{BU}(X)$ and $(bX,\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$. Then
(a’) $\mathcal{U}_{X}^{p}|_{\imath(G)}\in\mathbb{BU}(G)$, the completion of $(G,\mathcal{U}_{X}^{p}|_{\imath(G)})$ (the compactification of $G$) is the closure of $\imath(G)$ in $(bX)^{X}$,
(b’) for the $G$-Tychonoff space $(G=(G,\tau_{p}),bX,\tilde{\alpha})$ $\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)}\in\mathbb{E}(G)$, the completion of $(G,\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)})$ (the proper enveloping Ellis semigroup compactification of $G$) is the closure of $\tilde{\imath}(G)$ in $(bX)^{bX}$ ($\mbox{{\rm cl}}~\tilde{\imath}(G)\in\mathfrak{E}(G)$),
(c’) $\mathcal{U}_{X}^{p}|_{\imath(G)}\subset\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)}$, $(\mbox{{\rm cl}}~\imath(G),\imath)\leq(\mbox{{\rm cl}}~\tilde{\imath}(G),\tilde{\imath})$.
IV. Equiuniformity $\mathcal{U}_{X}^{p}|_{\imath(G)}$ (respectively, $\mathcal{U}_{\tilde{X}}^{p}|_{\tilde{\imath}(G)}$) is the least uniformity $\mathcal{U}$ on $G$ such that the orbit maps $\alpha_{x}:(G,\mathcal{U})\to(X,\mathcal{U}_{X})$, $x\in X$ (respectively, $\tilde{\alpha}_{x}:(G,\mathcal{U})\to(\tilde{X},\tilde{\mathcal{U}}_{X})$, $x\in\tilde{X}$) are uniformly continuous. Hence, $\mathcal{U}_{X}^{p}|_{\imath(G)}$ (respectively, $\mathcal{U}_{\tilde{X}}^{p}|_{\tilde{\imath}(G)}$) is the initial uniformity with respect to the correspondent family of maps.
Proof.
I. (1) is shown in § 1.2.
(2) From § 1.3 and Proposition 2.6 one has that $\imath$ is a topological isomorphism of $G$ onto the subsemigroup of $X^{X}$. The diagram
$\begin{array}[]{ccl}G\times X^{X}&\stackrel{{\scriptstyle\imath\times\mbox{{\rm id}}}}{{\longrightarrow}}&\imath(G)\times X^{X}\\ \quad\alpha_{\Delta X}\searrow&&\swarrow\circ\\ &X^{X}&\end{array}$
is commutative by item (2) of Proposition 2.6 and $\imath\times\mbox{{\rm id}}$ is a homeomorphism. Continuity of $\alpha_{\Delta X}$ yields continuity of $\circ$. (Continuity of the diagonal action $\alpha_{\Delta X}:G\times X^{X}\to X^{X}$ is equivalent to continuity of the multiplication $\circ:\imath(G)\times X^{X}\to X^{X}$.)
(3) follows from item (3) of Proposition 2.6.
(4) Since the diagonal action $\alpha_{\Delta X}$ is continuous and $\imath(G)$ is an invariant subset of $X^{X}$ the closure $\mbox{{\rm cl}}~\imath(G)$ is an invariant subset of $X^{X}$ and, hence, by Proposition 2.6,
$\imath(G)\circ\mbox{{\rm cl}}~\imath(G)=\bigcup\{\imath(g)\circ f\ |\ g\in G,f\in\mbox{{\rm cl}}~\imath(G)\}=\bigcup\{\alpha_{\Delta X}(g,f)\ |\ g\in G,f\in\mbox{{\rm cl}}~\imath(G)\}=\mbox{{\rm cl}}~\imath(G).$
Since $\imath(G)\circ f\subset\mbox{{\rm cl}}~\imath(G)$, $f\in\mbox{{\rm cl}}~\imath(G)$, and $X^{X}$ is a right topological semigroup, $\mbox{{\rm cl}}~\imath(G)\circ f\subset\mbox{{\rm cl}}~\imath(G)$, $f\in\mbox{{\rm cl}}~\imath(G)$. Hence,
$\mbox{{\rm cl}}~\imath(G)\circ\mbox{{\rm cl}}~\imath(G)=\mbox{{\rm cl}}~\imath(G)\ \mbox{and}\ \mbox{{\rm cl}}~\imath(G)\ \mbox{is a subsemigroup of}\ X^{X}.$
The second statement follows from item (2).
II (a), III (a’). The restriction $\mathcal{U}_{X}^{p}|_{\imath(G)}$ of the equiuniformity $\mathcal{U}_{X}^{p}$ on the invariant subset $\imath(G)$ is an equiuniformity on $\imath(G)$. If $\mathcal{U}_{X}\in\mathbb{BU}_{X}$, then $\mathcal{U}_{X}^{p}$ and its restriction $\mathcal{U}_{X}^{p}|_{\imath(G)}$ are totally bounded. The equivalence of $G$-spaces $(G,G,\cdot)$ and $(G,\imath(G),\alpha_{\Delta X}|_{G\times\imath(G)})$ and the fact that the completions of uniform subspaces are their closures in the completions of the correspondent uniform spaces are equal finishes the proof.
The proofs of II (b) and III (b’) are the same as of II (a) and III (a’), taking into consideration the fact that $\tilde{X}^{\tilde{X}}$ and $(bX)^{bX}$ are complete spaces, semigroups and by I (4) the closures of $G$ are invariant subsets.
III (c) follows from the uniform continuity of the projection of $(\tilde{X},\tilde{\mathcal{U}}_{X})^{\tilde{X}}$ onto $(\tilde{X},\tilde{\mathcal{U}}_{X})^{X}$, that $(X,\mathcal{U}_{X})^{X}$ is a uniform subspace of $(\tilde{X},\tilde{\mathcal{U}}_{X})^{X}$ and commutativity of the diagram
$\begin{array}[]{ccl}(\tilde{X},\tilde{\mathcal{U}}_{X})^{\tilde{X}}\stackrel{{\scriptstyle\mbox{{\rm pr}}}}{{\longrightarrow}}&(\tilde{X},\tilde{\mathcal{U}}_{X})^{X}&\hookleftarrow X^{X}\\ \imath^{\prime}\nwarrow&\tilde{\imath}\uparrow&\nearrow\imath\\ &G&\end{array}$
III (c’) is proved similarly. The inclusion $\mathcal{U}_{X}^{p}|_{\imath(G)}\subset\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)}$ yields the map of the corresponding compactifications.
IV. For arbitrary $x\in X$ and ${\rm U}\in\mathcal{U}_{X}$ take ${\rm V}=\{(g,h)\in G\times G\ |\ (\alpha(g,x),\alpha(h,x))\in{\rm U}\}\in\mathcal{U}_{X}^{p}|_{\imath(G)}$. If $(g,h)\in{\rm V}$, then $(\alpha(g,x),\alpha(h,x))\in{\rm U}$ and $\alpha_{x}:(G,\mathcal{U}_{X}^{p}|_{\imath(G)})\to(X,\mathcal{U}_{X})$ is uniformly continuous. The same argument works in the case of the completion $\tilde{X}$.
If $\mathcal{U}$ is an uniformity on $G$ and the orbit maps $\alpha_{x}:(G,\mathcal{U})\to(X,\mathcal{U}_{X})$, $x\in X$, are uniformly continuous, then the map $\imath:(G,\mathcal{U})\to(\imath(G),\mathcal{U}_{X}^{p}|_{\imath(G)})$ is uniformly continuous. Thus, $\mathcal{U}_{X}^{p}|_{\imath(G)}\subset\mathcal{U}$. The same argument works in the case of the completion $\tilde{X}$. ∎
(B) Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-space, $q_{\sigma}:G\to G/\mbox{{\rm St}}_{\sigma}$ is the quotient map, $\sigma\in\Sigma_{X}$.
The following diagram (the maps are continuous) is commutative for any $\sigma\in\Sigma_{X}$
[…]
If $(G=(G,\tau_{p}),X,\alpha)$ is a $G$-Tychonoff space and $\mathcal{U}_{X}$ is an equiuniformity on $X$, then the uniformity $\mathcal{U}_{X}^{p}|_{\imath(G)}$ on $\imath(G)$ has a spectral representation. It is initial with respect to the restrictions to $\imath(G)$ of the projections $\mbox{{\rm pr}}_{\sigma}:X^{X}\to(X^{\sigma},\mathcal{U}_{X}^{\sigma})$ (products $X^{\sigma}$ with the Cartesian product of equiuniformities $\mathcal{U}_{X}$).
$(X^{X},\mathcal{U}_{X}^{p})$ is uniformly equivalent to the inverse limit $\lim\limits_{\longleftarrow}S$ of the inverse spectrum $S=\{(X^{\sigma},\mathcal{U}_{X}^{\sigma}),\mbox{{\rm pr}}^{\sigma^{\prime}}_{\sigma},\Sigma_{X}\}$ (the proof is standard and the case of homeomorphism is in [14, Example 2.5.3]). Therefore, the completion of $(G,\mathcal{U}_{X}^{p}|_{\imath(G)})$ is uniformly equivalent to the limit of the inverse spectrum $\{\widetilde{\mbox{{\rm pr}}_{\sigma}(\imath(G))},\widetilde{\mbox{{\rm pr}}^{\sigma^{\prime}}_{\sigma}},\Sigma_{X}\}$, where $(\widetilde{\mbox{{\rm pr}}_{\sigma}(\imath(G))},\widetilde{\mathcal{U}}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$ is the completion of $(\mbox{{\rm pr}}_{\sigma}(\imath(G)),\mathcal{U}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$, $\widetilde{\mbox{{\rm pr}}^{\sigma^{\prime}}_{\sigma}}:\widetilde{\mbox{{\rm pr}}_{\sigma^{\prime}}(\imath(G))}\to\widetilde{\mbox{{\rm pr}}_{\sigma}(\imath(G))}$ is the extension of the restriction to $\mbox{{\rm pr}}_{\sigma^{\prime}}(\imath(G))$ of the projection $\mbox{{\rm pr}}^{\sigma^{\prime}}_{\sigma}:X^{\sigma^{\prime}}\to X^{\sigma}$, $\sigma\subset\sigma^{\prime}$.
The maximal equiuniformity $\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}}$ on $G/\mbox{{\rm St}}_{\sigma}$ is the quotient of the right uniformity $R$ on $G$ (follows from [37] and [11]).
Proposition 2.8.
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}$ is an equiuniformity on $X$. Then for any $\sigma\in\Sigma_{X}$ the map $\varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to((\mbox{{\rm pr}}_{\sigma}\circ\imath)(G),\mathcal{U}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$ is a uniformly continuous bijection. Hence,
$\mathcal{U}_{X}^{p}|_{\imath(G)}\subset R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G)).$
Proof.
Immediately follows from the commutativity of the diagram $(d)$, the condition that the restriction $\mathcal{U}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))}$ is an equiuniformity which is the restriction of the quotient uniformity of $\mathcal{U}_{X}^{p}$ and that the maximal equiuniformity on $G/\mbox{{\rm St}}_{\sigma}$ is the quotient of the right uniformity $R$ on $G$.
Since $R_{\Sigma_{X}}$ is initial with respect to the quotient maps $q_{\sigma}:G\to(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})$, see § 1.4, and $\mathcal{U}_{X}^{p}|_{\imath(G)}$ is initial with respect to the restrictions to $\imath(G)$ of the projections $\mbox{{\rm pr}}_{\sigma}:X^{X}\to(X^{\sigma},\mathcal{U}_{X}^{\sigma})$ one has $\mathcal{U}_{X}^{p}|_{\imath(G)}\subset R_{\Sigma_{X}}$. ∎
Corollary 2.9.
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}$ is an equiuniformity on $X$. If the uniformities on $G$ which are initial with respect to the quotient maps $q_{\sigma}:G\to(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})$, $\sigma\in\Sigma_{X}$, and the maps $\mbox{{\rm pr}}_{\sigma}|_{\imath(G)}\circ\imath:G\to(X^{\sigma},\mathcal{U}_{X}^{\sigma})$ correspondently, coincide (in particular,
[…]
then
(a) $\mathcal{U}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G))$,
(b) for any $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}$ (in particular, the maximal equiuniformity on $X$) ${\mathcal{U}^{\prime}}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G))$ (in particular, the map $\varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to(\mbox{{\rm pr}}_{\sigma}(\imath(G)),{\mathcal{U}^{\prime}}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$, $\sigma\in\Sigma_{X}$, is a uniform equivalence).
Remark 2.10.
Since the quotient maps $q_{\sigma}:G\to G/\mbox{{\rm St}}_{\sigma}$ are open, if the condition (op) of Corollary 2.9 is valid, then the restrictions of projections $\mbox{{\rm pr}}_{\sigma}|_{\imath(G)}$, $\sigma\in\Sigma_{X}$, are open.
(C) Let $(G,X,\curvearrowright)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}$ is an equiuniformity on $X$, $\mathcal{U}_{X}^{u}$ is a u.u.c. and $\tau_{u}$ is the induced t.u.c. on $X^{X}$. Evidently, ${\mathcal{U}}_{X}^{p}\subset\mathcal{U}_{X}^{u}$ and $\tau_{u}\geq\tau_{p}$.
It is easy to check that the diagonal action $\alpha_{\Delta X}:G\times(X^{X},\tau_{u})\to(X^{X},\tau_{u})$ is continuous, $\mathcal{U}_{X}^{u}$ is an equiuniformity on $X^{X}$. Its restriction $\mathcal{U}_{X}^{u}|_{\imath(G)}$ on the invariant subset $G=\imath(G)$ is an equiuniformity on $G$ and the induced topology $\tau_{u}$ is the t.u.c.
Lemma 2.11.
Let $((G,\tau),X,\curvearrowright)$ be a $G$-Tychonoff space. For every equiuniformity $\mathcal{U}_{X}$ on $X$
$\tau\geq\tau_{u}.$
If $\tau=\tau_{p}$, then $\tau=\tau_{u}$.
Proof.
The restriction of the diagonal action on the invariant subset $\imath(G)$ is continuous. Identifying $G=\imath(G)$ by Proposition 2.6 one has that the action $\cdot$ of $(G,\tau)$ on $(G,\tau_{u})$ is continuous. Hence, $\tau\geq\tau_{u}$.
Since $\tau_{u}\geq\tau_{p}$, if $\tau=\tau_{p}$, then $\tau_{u}=\tau$. ∎
Remark 2.12.
The u.u.c. $\mathcal{U}_{X}^{u}|_{\imath(G)}$ is a right uniformity $R$ on $G$ [37]. The base of $N_{G}(e)$ is formed by internals of the sets
$O_{{\rm U}}=\{g\in G\ |\ (x,\alpha(g,x))\in{\rm U},\ x\in X\},\ {\rm U}\in\mathcal{U}_{X}.$
$O^{-1}=O$, since $(\alpha(g^{-1},x),x)\in{\rm U}$, $x\in X$, iff $(y,\alpha(g,y))\in{\rm U}$, $y{\rm(}=\alpha(g^{-1},x){\rm)}\in X$.
Corollary 2.13.
Let $((G,\tau_{p}),X,\curvearrowright)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}$ is an equiuniformity on $X$. Then for any $g\in G$ and any ${\rm U}\in\mathcal{U}_{X}$ there exist $\sigma\in\Sigma_{X}$ and ${\rm V}\in\mathcal{U}_{X}$ such that
$\mbox{if}\ (\alpha(g,x),\alpha(h,x))\in{\rm V},\ x\in\sigma,\ \mbox{then}\ (\alpha(g,x),\alpha(h,x))\in{\rm U},\ x\in X.$
2.3. Ellis equiuniformity on a group from an equiuniformity on a phase space
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathbb{BU}(X)$ is a poset of totally bounded equiuniformities on $X$.
For any $\mathcal{U}_{X}\in\mathbb{BU}(X)$ by Theorem 2.7 (tem III (a’)) $\mathcal{U}_{X}^{p}|_{\imath(G)}\in\mathbb{BU}(G)$ and (item III (b’)) $\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)}\in\mathbb{E}(G)$ are defined. Thus, the maps
$\mathfrak{E}_{\alpha}(=\mathfrak{E}_{\alpha}(G,X,\alpha)):\mathbb{BU}(X)\to\mathbb{BU}(G),\ \mathfrak{E}_{\alpha}(\mathcal{U}_{X})=\mathcal{U}_{X}^{p}|_{\imath(G)}\ \mbox{and}$
$\tilde{\mathfrak{E}}_{\alpha}(=\tilde{\mathfrak{E}}_{\alpha}(G,X,\alpha)):\mathbb{BU}(X)\to\mathbb{E}(G),\ \tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=\mathcal{U}_{bX}^{p}|_{\tilde{\imath}(G)}$
are defined. The completion of $(G,\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}))$ is $(E(G,\mathcal{U}_{X}),\widetilde{\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})})$, where $E(G,\mathcal{U}_{X})=\mbox{{\rm cl}}~\tilde{\imath}(G)$ is the proper enveloping Ellis semigroup compactification.
Proposition 2.14.
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space. For any $\mathcal{U}_{X}\in\mathbb{BU}(X)$ $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$.
Let $(bX,\tilde{\mathcal{U}}_{X})$ be the completion of $(X,\mathcal{U}_{X})$. $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$ iff
()( ⋆ ) $\begin{array}[]{l}\mbox{for any}\ y\in bX\setminus X\ \mbox{and any}\ \tilde{\rm U}\in\tilde{\mathcal{U}}_{X}\ \mbox{there exists}\ \sigma\in\Sigma_{X}\ \mbox{and}\ \tilde{\rm V}\in\tilde{\mathcal{U}}_{X}\\ \mbox{such that if}\ (\tilde{\alpha}(g,x),\tilde{\alpha}(h,x))\in\tilde{\rm V},x\in\sigma,\ \mbox{then}\ (\tilde{\alpha}(g,y),\tilde{\alpha}(h,y))\in\tilde{\rm U}.\end{array}$
The maps $\mathfrak{E}_{\alpha}:\mathbb{BU}(X)\to\mathbb{BU}(G)$ and $\tilde{\mathfrak{E}}_{\alpha}:\mathbb{BU}(X)\to\mathbb{E}(G)$ preserve partial orders.
Proof.
$\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$ by item III (c’) of Theorem 2.7.
Condition ()($\star$)( ⋆ ) is a reformulation of the fact that the restriction of projection $\mbox{{\rm pr}}:(bX)^{bX}\to(bX)^{X}$ onto $\tilde{\imath}(G)$ is a uniform equivalence.
Let $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}\in\mathbb{BU}(X)$. For the $G$-compactifications $(bX=b_{\mathcal{U}_{X}}X,b=b_{\mathcal{U}_{X}})$ and $(b^{\prime}X=b_{\mathcal{U}^{\prime}_{X}}X,b^{\prime}=b_{\mathcal{U}^{\prime}_{X}})$ of $X$ (completions of $(X,\mathcal{U}_{X})$ and $(X,\mathcal{U}^{\prime}_{X})$ respectively) let a $G$-map $\varphi:b^{\prime}X\to bX$ be such that $\varphi\circ b^{\prime}=b$ (the map of $G$-compactifications). The restriction $\varphi|_{b^{\prime}(X)(=X)}:(X,\mathcal{U}^{\prime}_{X})\to(X,\mathcal{U}_{X})$ is uniformly continuous.
The Cartesian product $\varphi^{X}:b^{\prime}X^{X}\to bX^{X}$ of maps $\varphi$ of compacta is uniformly continuous. Hence, the restriction $\varphi^{X}|_{\imath^{\prime}(G)(=G)}:(G,\mathfrak{E}_{\alpha}(U^{\prime}_{X}))\to(G,\mathfrak{E}_{\alpha}(U_{X}))$ is uniformly continuous and $\mathfrak{E}_{\alpha}(U_{X})\subset\mathfrak{E}_{\alpha}(U^{\prime}_{X})$.
Let $\tilde{\alpha}:G\times bX\to bX$ and $\tilde{\alpha}^{\prime}:G\times b^{\prime}X\to b^{\prime}X$ be the extensions of $\alpha:G\times X\to X$, $X^{\prime}\subset b^{\prime}X$ is any subset the restriction of $\varphi$ to which is a bijection. The Cartesian product $\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{b^{\prime}X}$ of maps $\varphi$ is a $G$-map. The projection $\mbox{{\rm pr}}_{X^{\prime}}:(bX)^{b^{\prime}X}\to(bX)^{X^{\prime}}$ is a $G$-map, the map $\jmath:(bX)^{X^{\prime}}\to(bX)^{bX}$, $\jmath(f)=h$, where $h(t)=f({\varphi^{-1}(t)})$, $t\in bX$ (identification of coordinates) is also a $G$-map. Their composition $\psi=\jmath\circ\mbox{{\rm pr}}_{X^{\prime}}\circ\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{bX}$ is a $G$-map.
The following diagram
$\begin{array}[]{ccl}(b^{\prime}X)^{b^{\prime}X}&\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}&(bX)^{bX}\\ \imath^{\prime}\nwarrow&&\nearrow\imath\\ &G&\end{array}$
is commutative. In fact, for any $x\in X$
$\imath(g)(b(x))=\tilde{\alpha}(g,b(x))=\alpha(g,b(x))\ \mbox{and}$
$[(\psi\circ\imath^{\prime})(g)](b(x))=\mbox{{\rm pr}}_{b(x)}\big{(}(\psi\circ\imath^{\prime})(g)\big{)}=\mbox{{\rm pr}}_{b(x)}\big{(}(\jmath\circ\mbox{{\rm pr}}_{X^{\prime}}\circ\varphi^{b^{\prime}X}\circ\imath^{\prime})(g)\big{)}=$
since $\mbox{{\rm pr}}_{b(x)}\circ\jmath\circ\mbox{{\rm pr}}_{X^{\prime}}=\mbox{{\rm pr}}_{b^{\prime}(x)}$
$=\mbox{{\rm pr}}_{b^{\prime}(x)}\big{(}(\varphi^{b^{\prime}X}\circ\imath^{\prime})(g)\big{)}=\varphi\big{(}\imath^{\prime}(g)(b^{\prime}(x))\big{)}=\varphi\big{(}\tilde{\alpha}^{\prime}(g,b^{\prime}(x))\big{)}=$
since $\varphi$ is a $G$-map
$=\tilde{\alpha}(g,\varphi(b^{\prime}(x)))=\tilde{\alpha}(g,b(x))\stackrel{{\scriptstyle x\in X}}{{=}}\alpha(g,b(x)).$
Thus, $\imath(g)(t)=(\psi\circ\imath^{\prime})(g)(t)$, $t\in b(X)$. Being continuous they coincide on $bX$ and $\imath(g)=(\psi\circ\imath^{\prime})(g)$.
Continuity of $\psi$, commutativity of the above diagram and compactness (by Theorem 2.7) of $\mbox{{\rm cl}}~\imath^{\prime}(G)=E(G,\mathcal{U}^{\prime}_{X})$, $\mbox{{\rm cl}}~\imath(G)=E(G,\mathcal{U}_{X})$ yields that the restriction $\psi|_{E(G,\mathcal{U}^{\prime}_{X})}:E(G,\mathcal{U}^{\prime}_{X})\to E(G,\mathcal{U}_{X})$ is the map of $G$-compactifications. From the proof of Proposition 2.5 one has $E(G,\mathcal{U}^{\prime}_{X})\geq E(G,\mathcal{U}_{X})$. It remains to apply Proposition 2.5 to obtain $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}^{\prime}_{X})$. ∎
Remark 2.15.
The map $\psi$ of proper Ellis semigroup compactifications in Proposition 2.14 doesn’t depend on the choice of $X^{\prime}$ in Proposition 2.14.
In general, the maps $\mathfrak{E}_{\alpha}:\mathbb{BU}(X)\to\mathbb{BU}(G)$ and $\tilde{\mathfrak{E}}_{\alpha}:\mathbb{BU}(X)\to\mathbb{E}(G)$ are neither injective, no surjective.
If $X$ is compact, then $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$ for the unique uniformity $\mathcal{U}_{X}$ on $X$.
Proposition 2.16.
([42, Theorem 2.1], [41, Theorem 1]) Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space and $\mathcal{U}_{X}\in\mathbb{BU}(X)$.
If for any $\sigma\in\Sigma_{X}$ and for any entourage ${\rm U}\in{\mathcal{U}}_{X}$ there is an entourage ${\rm V}={\rm V}(\sigma;{\rm U})\in{\mathcal{U}}_{X}$ such that the condition is met:
()( ⋆ ⋆ ) $\begin{array}[]{c}\mbox{\rm if for}\ g,h\in G\ (\alpha(g,x),\alpha(h,x))\in{\rm V}\ \mbox{\rm holds for}\ x\in\sigma,\\ \mbox{\rm then there exists}\ \ g^{\prime}\in g\mbox{{\rm St}}_{\sigma}\ \ \mbox{\rm such that}\ \ h\in O_{{\rm U}}g^{\prime},\\ \end{array}$
where $O_{{\rm U}}=\{f\in G\ |\ (\alpha(f,x),x)\in{\rm U},\ x\in X\}$. Then $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{X}}$.
Proof.
By Corollary 2.9 it is sufficient to show that for any $\sigma\in\Sigma_{X}$ the map $\varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to((\mbox{{\rm pr}}_{\sigma}\circ\imath)(G),{\mathcal{U}}_{X}^{\sigma}|_{(\mbox{{\rm pr}}_{\sigma}\circ\imath)(G)})$ is a uniform equivalence. Since $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\subset R_{\Sigma_{X}}$ by Lemma 2.8, only the uniform continuity of $\varphi_{\sigma}^{-1}$ must be checked, i.e. in the cover
$\{O_{{\rm U}}g\mbox{{\rm St}}_{\sigma}\ |\ g\in G\},\ {\rm U}\in{\mathcal{U}}_{X},$
the cover
$\{U_{\sigma;g;{\rm V}}=\{h\in G\ |\ (\alpha(g,x),\alpha(h,x))\in{\rm V},\ x\in\sigma\}\ |\ g\in G\},$
is refined for some ${\rm V}\in{\mathcal{U}}_{X}$.
[…]
$h\in O_{{\rm U}}g\mbox{{\rm St}}_{\sigma}.$
Therefore, $U_{\sigma;g;{\rm V}}\subset O_{{\rm U}}g\mbox{{\rm St}}_{\sigma}$, i.e. the cover $\{U_{\sigma;g;{\rm V}}\ |\ g\in G\}\succ\{O_{{\rm U}}g\mbox{{\rm St}}_{\sigma}\ |\ g\in G\}$ and, hence, taking in account item (C) of § 2.2, $\varphi_{\sigma}$ is a uniform equivalence. ∎
Remark 2.17.
The condition $h\in O_{{\rm U}}g$ in Proposition 2.16 is equivalent to the condition $(\alpha(h,x),\alpha(g,x))\in{\rm U}$, $x\in X$.
Indeed, $(\alpha(h,x),\alpha(g,x))\in{\rm U},x\in X\Longleftrightarrow(\alpha(hg^{-1},x),x)\in{\rm U},x\in X\Longleftrightarrow hg^{-1}\in O_{{\rm U}}\Longleftrightarrow h\in O_{{\rm U}}g$.
Corollary 2.18.
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}\in\mathbb{BU}(X)$ and $(bX,\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$ ($G$-compactification of $X$). Then
  • (1)
    $\begin{array}[]{ccl}\mathfrak{E}_{\alpha}(\mathcal{U}_{X})&\subset&R_{\Sigma_{X}}\\ \cap&&\cap\\ \tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})&\subset&R_{\Sigma_{bX}}.\end{array}$
  • (2)
[…]
  • (3)
[…]
Remark 2.19.
[…]
Remark 2.20.
(1) The above approach of obtaining compactifications of transformation groups is used in [41] for the group of homeomorphisms of the compactum $K$ in the topology of pointwise convergence.
[…]
Other results that guarantee the Roelcke precompactness of an acting group are the following.
(a) If a (closed) subgroup $H$ of a topological group $G$ is Roelcke precompact and the maximal equiuniformity on the coset space $G/H$ (with the action of $G$ by multiplication on the left) is totally bounded, then $G$ is Roelcke precompact [20, Proposition 6.4] ($G$ can be considered as a transformation group of $G/H$).
(b) If a topological group $G$ has a directed family $\mathcal{K}$ of small subgroups such that the maximal equiuniformity on a coset space $G/H$ (with the action of $G$ by multiplication on the left) is totally bounded, $H\in\mathcal{K}$, then $G$ is Roelcke precompact [26, Corollary 4.5].
(c) If the action $G\curvearrowright X$, where $X$ is a discrete space, is oligomorphic, then $(G,\tau_{\partial})$ is Roelcke precompact [44] (equivalently, for the action of $G$ in the permutation topology (which coincides with the topology of pointwise convergence) the maximal equiuniformity on $X$ is totally bounded [42, Theorem 3.3]).
Proposition 2.21.
Let $(G=(G,\tau_{p}),X,\alpha)$ and $(G=(G,\tau_{p}),Y,\alpha^{\prime})$ be $G$-Tychonoff spaces, $\mathcal{U}_{X}\in\mathbb{BU}(X)$ and $(bX,\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$ ($G$-compactification of $X$), $\mathcal{U}_{Y}\in\mathbb{BU}(Y)$ and $(bY,\tilde{\mathcal{U}}_{Y})$ is the completion of $(Y,\mathcal{U}_{Y})$ ($G$-compactification of $Y$), $\varphi:(X,\mathcal{U}_{X})\to(Y,\mathcal{U}_{Y})$ is a uniformly continuous $G$-map, $\varphi(X)$ is dense in $Y$. Then
$\tilde{\mathfrak{E}}_{\alpha^{\prime}}(\mathcal{U}_{Y})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}),\ R_{\Sigma_{bY}}\subset R_{\Sigma_{bX}}.$
If $\varphi$ is a surjection, then
$\mathfrak{E}_{\alpha^{\prime}}(\mathcal{U}_{Y})\subset\mathfrak{E}_{\alpha}(\mathcal{U}_{X}),\ R_{\Sigma_{Y}}\subset R_{\Sigma_{X}}.$
Proof.
Let $\tilde{\varphi}:bX\to bY$ be the extension (surjection) of $\varphi$ (a uniformly continuous surjective $G$-map). For any $y\in bY$ and $\tilde{\varphi}(x)=y$ the following is valid $\tilde{\alpha}^{\prime}_{y}=\tilde{\varphi}\circ\tilde{\alpha}_{x}$.
By Theorem 2.7 item IV $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\in\mathbb{E}(G)$ us the least equiuniformity $\mathcal{U}$ on $G$ such that the orbit maps $\tilde{\alpha}_{x}:(G,\mathcal{U})\to(bX,\tilde{\mathcal{U}}_{X})$, $x\in bX$, are uniformly continuous. Hence, the orbit maps $\tilde{\alpha}^{\prime}_{y}:(G,\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}))\to(bY,\tilde{\mathcal{U}}_{Y})$, $y\in bY$, are uniformly continuous. Again applying Theorem 2.7 item IV, one obtains $\tilde{\mathfrak{E}}_{\alpha^{\prime}}(\mathcal{U}_{Y})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$.
Since $\tilde{\varphi}$ is a $G$-map $\mbox{{\rm St}}_{x}\subset\mbox{{\rm St}}_{\varphi(x)}$ (if $\alpha(g,x)=x$, then $\alpha^{\prime}(g,\varphi(x))=\varphi(\alpha(g,x))=\varphi(x)$). The usage of definitions of $R_{\Sigma_{bY}}$ and $R_{\Sigma_{bX}}$ finishes the proof of the inclusion $R_{\Sigma_{bY}}\subset R_{\Sigma_{bX}}$.
The proof of the second statement is analogous. ∎
2.4. The maps $\mathfrak{E}_{*}$ and $\tilde{\mathfrak{E}}_{*}$
For a $G$-Tychonoff space $(G,G,\cdot)$ the t.p.c. $\tau_{p}$ coincides with the topology of $G$ by Lemma 1.3 and, hence, from the considerations in § 2.3 the maps $\mathfrak{E}_{*}:\mathbb{BU}(G)\to\mathbb{BU}(G)$, $\tilde{\mathfrak{E}}_{*}:\mathbb{BU}(G)\to\mathbb{E}(G)$ are correctly defined.
Theorem 2.22.
For a topological group $G$ and $\mathcal{U},\mathcal{U}^{\prime}\in\mathbb{BU}(G)$ the following hold.
  • (a)
    $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$.
  • (b)
    $\mathcal{U}=\mathfrak{E}_{*}(\mathcal{U})$ iff for any $x\in G$ the orbit map $\cdot_{x}:(G,\mathcal{U})\to(G,\mathcal{U})$, $\cdot_{x}(g)=gx$, is uniformly continuous (in particular, $\mathfrak{E}_{*}(\mathfrak{E}_{*}(\mathcal{U}))=\mathfrak{E}_{*}(\mathcal{U})$).
  • (c)
    If $\mathcal{U}\subset\mathcal{U}^{\prime}$, then $\mathfrak{E}_{*}(\mathcal{U})\subset\mathfrak{E}_{*}(\mathcal{U}^{\prime})$ (in particular, if $\mathcal{U}\subset\mathcal{U}^{\prime}\subset\mathfrak{E}_{*}(\mathcal{U})$, then $\mathfrak{E}_{*}(\mathcal{U})=\mathfrak{E}_{*}(\mathcal{U}^{\prime})$).
  • (d)
    If $\mathcal{U}\in\mathbb{E}(G)\subset\mathbb{BU}(G)$, then $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\mathfrak{E}_{*}(\mathcal{U})=\mathcal{U}$ (in particular, $\tilde{\mathfrak{E}}_{*}:\mathbb{BU}(G)\to\mathbb{E}(G)$ is a surjection, $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}(\tilde{\mathfrak{E}}_{*}(\mathcal{U}))$ and $\mathcal{U}\in\mathbb{E}(G)$ iff $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\mathcal{U}$).
  • (e)
    $\mathcal{U}\in\mathbb{E}(G)$ iff for any $x\in bG$ ($(bG,\tilde{\mathcal{U}})$ is the completion of $(G,\mathcal{U})$) the orbit map ~$\tilde{\cdot}_{x}:(G,\mathcal{U})\to(bG,\tilde{\mathcal{U}})$, ~$\tilde{\cdot}_{x}(g)=\tilde{\cdot}(g,x)$, is uniformly continuous.
  • (f)
    If $\mathcal{U}\subset\mathcal{U}^{\prime}$, then $\tilde{\mathfrak{E}}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})$ (in particular, if $\mathcal{U}\subset\mathcal{U}^{\prime}\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$, then $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})$ and $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}({\mathfrak{E}}_{*}(\mathcal{U}))$).
Proof.
(a) The projection $\mbox{{\rm pr}}_{e}:(G^{G},\mathcal{U}^{p})\to(G=\{e\}\times G,\mathcal{U})$ is uniformly continuous. Therefore, $\mbox{{\rm pr}}_{e}|_{\imath(G)}:(\imath(G),\mathfrak{E}_{*}(\mathcal{U}))\to(G,\mathcal{U})$ is uniformly continuous. Hence, $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})$. The inclusion $\mathfrak{E}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ is proved in Proposition 2.14.
(b) follows from the inclusion $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})$ in (a) and Theorem 2.7 item IV.
(c) The preservation of order follows from Proposition 2.14. The rest follows from the inclusion $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})$ and Theorem 2.7 item IV.
(d) By (a) $\mathcal{U}\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$.
Let $(bG,b)$ be the $G$-compactification of $G$ ($(bG,\tilde{\mathcal{U}})$ is the completion of $(G,\mathcal{U})$), ~$\tilde{\cdot}:G\times bG\to bG$ is the extension of action $\cdot$, $\tilde{\imath}:G\to(bG)^{bG}$.
Since $(bG,\bullet)$ is a right topological semigroup and the restriction of multiplication on $b(G)\times bG$ is continuous:
$\imath(g)(t)=\tilde{\cdot}(g,t)=b(g)\bullet t,\ t\in bG,\ \mbox{and}$
the continuous multiplication on the right is uniformly continuous. Therefore, for any $t\in bG$ and for any ${\rm U}\in\mathcal{U}$ there exists ${\rm V}\in\mathcal{U}$ such that
$\mbox{if for}\ f,g\in G,\ (f,g)\in{\rm V},\ \mbox{then}\ (b(f)\bullet t,b(g)\bullet t)\in{\rm U}.$
Therefore, for any entourage $\hat{\rm U}=\{(f,g)\in G\times G\ |\ (\imath(f)(t)=b(f)\bullet t,\imath(g)(t)=b(g)\bullet t)\in{\rm U}\}$, where $t\in bG$, ${\rm U}\in\mathcal{U}$, from the subbase of $\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ there exists ${\rm V}\in\mathcal{U}$ such that
$\mbox{if}\ (f,g)\in{\rm V},\ \mbox{then}\ (f,g)\in\hat{\rm U}$
and the map $\tilde{\imath}:G\to\tilde{\imath}(G)$ is uniformly continuous. Hence, $\tilde{\mathfrak{E}}_{*}(\mathcal{U})\subset\mathcal{U}$ and, finally $\mathcal{U}=\tilde{\mathfrak{E}}_{*}(\mathcal{U})$.
The inclusion $\mathbb{E}(G)\subset\mathbb{BU}(G)$ yields that $\tilde{\mathfrak{E}}_{*}$ is surjective. Idempotentness of $\tilde{\mathfrak{E}}_{*}$ is evident.
[…]
(f) The first statement follows from Proposition 2.14. If $\mathcal{U}\subset\mathcal{U}^{\prime}\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$, then
$\tilde{\mathfrak{E}}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})\subset\tilde{\mathfrak{E}}_{*}(\tilde{\mathfrak{E}}_{*}(\mathcal{U}))\stackrel{{\scriptstyle(d)}}{{=}}\tilde{\mathfrak{E}}_{*}(\mathcal{U})$
and $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})$. Applying (a) the last statement is obtained. ∎
Example 2.23.
1. The greatest ambit of $G$ is the pair $(\beta_{G}G,e)$ where $\beta_{G}G$ is the completion of $G$ with respect to the maximal totally bounded equiuniformity $\mathcal{U}_{{\rm max}}$ on $G$ [7].
$\mathcal{U}_{{\rm max}}\in\mathbb{E}(G)$, the greatest ambit $\beta_{G}G$ is the maximal proper enveloping Ellis semigroup compactification of $G$.
Indeed, $\mathcal{U}_{{\rm max}}\subset\mathfrak{E}_{*}(\mathcal{U}_{{\rm max}})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U}_{{\rm max}})\subset\mathcal{U}_{{\rm max}}\in\mathbb{BU}(G)$ by Theorem 2.22 and the maximality of $\mathcal{U}_{{\rm max}}\in\mathbb{BU}(G)$. Hence, $\mathcal{U}_{{\rm max}}=\tilde{\mathfrak{E}}_{*}(\mathcal{U}_{{\rm max}})\in\mathbb{E}(G)$.
If $R=\mathcal{U}_{{\rm max}}$, then $G$ is a precompact group (see, for example, [3, § 3.7]) and the unique (see, for example, [25]) proper enveloping Ellis semigroup compactification of $G$ is a topological group.
For a non-Archimedian group $G$ the spectral description of $\beta_{G}G$ is given in [36, Corollary 3.3].
2. Let $\mathcal{U}_{L\wedge R}$ be the precompact reflection of the Roelcke uniformity $L\wedge R$. $\mathcal{U}_{L\wedge R}={\mathfrak{E}}_{*}(\mathcal{U}_{L\wedge R})$.
Indeed, for any cover $\{OgO\ |\ g\in G\}\in L\wedge R$, $O\in N_{G}(e)$, and any $h\in G$ there exists $V\in N_{G}(e)$ such that $Vh\subset hO$. Hence, the orbit map $\cdot_{h}:(G,L\wedge R)\to(G,L\wedge R)$, $\cdot_{h}(g)=gh$ is uniformly continuous ($\{VgV\ |\ g\in G\}h=\{VgVh\ |\ g\in G\}\succ\{OghO\ |\ g\in G\}=\{OgO\ |\ g\in G\}$). Therefore, for the precompact reflection $\mathcal{U}_{L\wedge R}$ of $L\wedge R$ the orbit map $\cdot_{h}:(G,\mathcal{U}_{L\wedge R})\to(G,\mathcal{U}_{L\wedge R})$, $\cdot_{h}(g)=gh$ is uniformly continuous
Since $L\wedge R$ is an equiuniformity for both actions: multiplication on the left and the right, its precompact reflection $\mathcal{U}_{L\wedge R}$ is a totally bounded equiuniformity for the multiplication on the left and the right. Hence, the actions $\cdot_{l}:G\times G\to G$, $\cdot_{l}(g,x)=gx$ and $\cdot_{r}:G\times G\to G$, $\cdot_{r}(g,x)=xg$ extend to the continuous actions on the completion of $(G,\mathcal{U}_{L\wedge R})$.
[…]
From Theorem 2.22 one has $\mathcal{U}_{L\wedge R}\in\mathbb{E}(G)$ iff $\mathcal{U}_{L\wedge R}=\tilde{\mathfrak{E}}_{*}(\mathcal{U}_{L\wedge R})$.
3. If $G$ is a locally compact group then, the least totally bounded uniformity $\mathcal{U}_{\alpha}$ on $G$ is an Ellis equiuniformity and the one-point Alexandroff compactification $\alpha G$ is a proper semitopological semigroup compactification of $G$ [5].
The action $\cdot:G\times G\to G$ extends to a continuous action ~$\tilde{\cdot}:G\times\alpha G\to\alpha G$. By Theorem 2.22 item (e) if for any $x\in\alpha G$ the map ~$\tilde{\cdot}_{x}:(G,\mathcal{U}_{\alpha})\to(\alpha G=G\cup\{\infty\},\tilde{\mathcal{U}}_{\alpha})$, ~$\tilde{\cdot}_{x}(g)=\tilde{\cdot}(g,x)$, is uniformly continuous, then $\mathcal{U}_{\alpha}\in\mathbb{E}(G)$ and $\alpha G\in\mathfrak{E}(G)$.
If $x\in G$, then the extension ~$\widetilde{\tilde{\cdot}_{x}}:\alpha G\to\alpha G$, ~$\widetilde{\tilde{\cdot}_{x}}(\infty)=\infty$, of the map ~$\tilde{\cdot}_{x}:G\to\alpha G$ is a continuous map and, hence, uniformly continuous. Therefore its restriction to $G$ is uniformly continuous.
If $x=\infty$, then ~$\tilde{\cdot}(g,\infty)=\infty$, the map ~$\tilde{\cdot}_{\infty}:G\to\alpha G$ is constant and, hence, uniformly continuous.
The extension $\tilde{\imath}:\alpha G\to\tilde{\imath}(\alpha G)\subset(\alpha G)^{\alpha G}$ of $\imath$ is a topological isomorphism of semigroups and $\tilde{\imath}(\infty)(x)=\widetilde{\tilde{\cdot}_{x}}(\infty)=\infty$, $x\in\alpha G$. Hence, $\tilde{\imath}(\infty)$ is a continuous map, the multiplication on the left in $\tilde{\imath}(\alpha G)$ is continuous, and $\alpha G$ is a proper semitopological semigroup compactification of $G$.
From Proposition 2.21, Theorem 2.7 item IV and Theorem 2.22 item (a) one has.
Corollary 2.24.
Let $(G=(G,\tau_{p}),X,\alpha)$ be a $G$-Tychonoff space, $\mathcal{U}_{X}\in\mathbb{BU}(X)$ and $(bX,\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$ ($G$-compactification of $X$), $\mathcal{U}\in\mathbb{BU}(G)$ and $(bG,\tilde{\mathcal{U}})$ is the completion of $(G,\mathcal{U})$, $\alpha_{x}:(G,\mathcal{U})\to(X,\mathcal{U}_{X})$ is a uniformly continuous $G$-map, $x\in X$, $\alpha_{x}(G)$ is dense in $X$ for some $x\in X$. Then $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\subset\tilde{\mathfrak{E}}_{\star}(\mathcal{U})$ and $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\subset\mathcal{U}\subset\mathfrak{E}_{\star}(\mathcal{U})$.
3. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for a uniformly equicontinuous action
An action $\alpha:G\times(X,\mathcal{L})\to(X,\mathcal{L})$ is uniformly equicontinuous if $\{\alpha^{g}:X\to X,\ \alpha^{g}(x)=\alpha(g,x)\ |\ g\in G\}$ is a uniformly equicontinuous family of maps on $X$. Equivalently, for any $u\in\mathcal{L}$ there exists $v\in\mathcal{L}$ such that $gv\succ u$, for any $g\in G$. If an action $G\curvearrowright(X,\mathcal{L})$ is uniformly equicontinuous, then the t.p.c. $\tau_{p}$ is the smallest admissible group topology on $G$ [26, Lemma 3.1]. Moreover, $((G,\tau_{p}),X,\alpha)$ is a $G$-Tychonoff space [29].
Remark 3.1.
Let $((G,\tau_{p}),X,\curvearrowright)$ be a $G$-space. If $X$ is a compactum and the action $G\curvearrowright X$ is (uniformly) equicontinuous, then the completion of $(G,{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}))$ is a compact topological group [47, Theorem 4.2] (see, also [9, Ch.10, § 3] and [26, Theorem 3.33]).
Let $H$ be a neutral subgroup of a topological group $G$ (see, [37, Definition 5.29]) without invariant subgroups. Then $G$ is a transitive uniformly equicontinuous group of homeomorphisms of $(G/H,\mathcal{L})$, where $\mathcal{L}$ is the quotient uniformity on $G/H$ of the left uniformity $L$ on $G$ (the final uniformity with respect to the map $(G,L)\to G/H$). By [37, Theorem 5.28] $H$ is a neutral subgroup of $(G,\tau_{p})$ and $((G,\tau_{p}),G/H,\alpha)$ is a $G$-Tychonoff space (with an open action[26, Proposition 3.2].
Theorem 3.2.
Let $H$ be a neutral subgroup without nontrivial invariant subgroups of a topological group $G$, $\mathcal{U}_{G/H}$ is the maximal equiuniformity on $G/H$ for the action $\alpha:(G,\tau_{p})\times G/H\to G/H$. If $\mathcal{U}_{G/H}\in\mathbb{BU}(G/H)$, then
(1)( 1 ) $\begin{array}[]{ccclc}{\mathfrak{E}}_{\alpha}(\mathcal{U}_{G/H})&\subset&L\wedge R&\subset&R_{\Sigma_{G/H}}\\ \cap&&&&\cap\\ \widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})&&\subset&&R_{\Sigma_{\beta_{G}{G/H}}},\end{array}$
(2) $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})\subset\widetilde{{\mathfrak{E}}}_{\star}(\mathcal{U}_{L\wedge R})$, where $\mathcal{U}_{L\wedge R}$ is the precompact reflection of $L\wedge R$,
(3) if $L\wedge R\in\mathbb{BU}(G)$ and $L\wedge R\subset\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})$, then $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})=\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)$,
[…]
Proof.
(1) The inclusions ${\mathfrak{E}}_{\alpha}(\mathcal{U}_{G/H})\subset\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})$, $R_{\Sigma_{G/H}}\subset R_{\Sigma_{\beta_{G}{G/H}}}$ and $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})\subset R_{\Sigma_{\beta_{G}{G/H}}}$ follow from Corollary 2.18. The inclusion $L\wedge R\subset R_{\Sigma_{G/H}}$ follows from Remark 1.5.
By Proposition 3.25 [26] $\mathcal{U}_{G/H}=\mathcal{L}\wedge\mathcal{R}$, where $\mathcal{L}\wedge\mathcal{R}$ is the quotient uniformity of $L\wedge R$ on $G$ for the quotient map $q:G\to G/H$ ($\mathcal{U}_{G/H}$ is the quotient uniformity of the right uniformity $R$ on $G$).
The inclusion ${\mathfrak{E}}_{\alpha}(\mathcal{U}_{G/H})\subset L\wedge R$ follows from the uniform continuity of the surjective orbit maps $\alpha_{x}:(G,L\wedge R)\to(G/H,\mathcal{L}\wedge\mathcal{R})$, $x\in G/H$, by Corollary 2.24.
(2) The surjective $G$-map $q:(G,L\wedge R)\to(G/H,\mathcal{L}\wedge\mathcal{R})$ is uniformly continuous. Since $\mathcal{L}\wedge\mathcal{R}\in\mathbb{BU}({G/H})$, for the precompact reflection $\mathcal{U}_{L\wedge R}$ of $L\wedge R$ the map $q:(G,\mathcal{U}_{L\wedge R})\to(G/H,\mathcal{L}\wedge\mathcal{R})$ is uniformly continuous. By Proposition 2.21 $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})\subset\widetilde{{\mathfrak{E}}}_{\star}(\mathcal{U}_{L\wedge R})$.
(3) $L\wedge R=\mathcal{U}_{L\wedge R}$. From the inclusions $L\wedge R\subset\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})\subset\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)$ (item (2)) by items (d) and (f) of Theorem 2.22 $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})=\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)$, since $\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})\in\mathbb{E}(G)$.
[…]
3.1. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an action on a discrete space
Let $G$ be a subgroup of the permutation group ${\rm S}(X)$ of a discrete (infinite) space $X$. A discrete space $X$ can be considered as a metric space with distance 111 between distinct points. Then ${\rm S}(X)$ is an isometry group and $G$ is its subgroup. The permutation topology $\tau_{\partial}$ on $G$ is the smallest admissible group topology (and coincides with the topology of pointwise convergence) for the action $\alpha:G\times X\to X$. The base of the nbds of the unit of $(G,\tau_{\partial})$ is formed by clopen subgroups
$\mbox{{\rm St}}_{\sigma}=\{g\in G\ |\ \alpha(g,x)=x,\ x\in\sigma\},\ \sigma\in\Sigma_{X},$
and $(G,\tau_{\partial})$ is non-Archimedean.
Lemma 3.3.
Let $X$ be discrete, $G$ is a subgroup of ${\rm S}(X)$. Then for $(G,\tau_{\partial})$
$L\wedge R=R_{\Sigma_{X}}.$
Proof.
The base of the Roelcke uniformity $L\wedge R$ on $(G,\tau_{\partial})$ is formed by the covers
$\{\mbox{{\rm St}}_{\sigma}g\mbox{{\rm St}}_{\sigma}\ |\ g\in G\},\ \sigma\in\Sigma_{X},$
which form the base of the uniformity $R_{\Sigma_{X}}$ defined by the family of stabilizers $\mbox{{\rm St}}_{\sigma}$, $\sigma\in\Sigma_{X}$ (see § 1.4). Hence, $L\wedge R=R_{\Sigma_{X}}$. ∎
3.1.1. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultratransitive action on a discrete space
Theorem 3.4.
Let $X$ be a discrete space and $\mathcal{U}_{X}$ is the maximal equiuniformity on $X$ for the action $\alpha:({\rm S}(X),\tau_{\partial})\times X\to X$. Then
  • (a)
    $\mathbb{BU}(X)=\{\mathcal{U}_{X}\}$ and $\tilde{X}=\alpha X$, where $(\tilde{X},\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$, ${}\alpha X=X\cup\{\infty\}$ is the Alexandroff one-point compactification of $X$,
  • (b)
    all inclusions in Theorem 3.2 item (1) are equalities and $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=L\wedge R$ (hence, $({\rm S}(X),\tau_{\partial})$ is Roelcke precompact and the Roelcke compactification $(b_{r}{\rm S}(X),b_{r})$ of $({\rm S}(X),\tau_{\partial})$ is a proper enveloping Ellis semigroup compactification),
  • (c)
    $b_{r}{\rm S}(X)$ is the set of selfmaps $f$ of $\alpha X$ in the t.p.c. such that $f(\infty)=\infty$, $f$ is a bijection on $Y\subset X$ and $f(X\setminus Y)=\infty$, where $Y$ is an arbitrary subset of $X$,
  • (d)
    $b_{r}{\rm S}(X)$ is a semitopological semigroup.
Proof.
(a) The base of the maximal equiuniformity $\mathcal{U}_{X}$ on $X$ is formed by the covers
$u_{\theta}=\{\mbox{{\rm St}}_{\theta}x=\alpha(\mbox{{\rm St}}_{\theta},x)\ |\ x\in X\},\ \theta\in\Sigma_{X},\ {\rm(}u_{\theta^{\prime}}\succ u_{\theta}\ \mbox{if}\ \theta\subset\theta^{\prime}{\rm)},$
see, for example, [11]. Since $\mbox{{\rm St}}_{\theta}x=x$ if $x\in\theta$ and $\mbox{{\rm St}}_{\theta}x=\alpha X\setminus\theta$ if $x\notin\theta$ they are of the form
$\{\{\{x\}\ |\ {x\in\theta}\},\ X\setminus\theta\},\ \theta\in\Sigma_{X},$
and coincides with the restriction on $X$ of the unique uniformity on $\alpha X$. Hence $\alpha X=\tilde{X}$ and the maximal equiuniformity $\mathcal{U}_{X}$ on $X$ is totally bounded and unique.
(b) From Lemma 3.3 $L\wedge R=R_{\Sigma_{X}}$. Since $\mbox{{\rm St}}_{\infty}={\rm S}(X)$ for the extended action $\tilde{\alpha}:{\rm S}(X)\times\alpha X\to\alpha X$ (continuity of each $g\in{\rm S}(X)$ yields that $\tilde{\alpha}(g,\infty)=\infty$), $R_{\Sigma_{X}}=R_{\Sigma_{\alpha X}}$. Since $\tilde{\alpha}_{\infty}(g)=\tilde{\alpha}(g,\infty)=\infty$, $g\in{\rm S}(X)$, the restriction to $\tilde{\imath}({\rm S}(X))$ of the projection $\mbox{{\rm pr}}_{X}:(\alpha X)^{\alpha X}\to(\alpha X)^{X}$ is a topological isomorphism and $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$.
[…]
$\begin{array}[]{c}\mbox{\rm if for}\ g,h\in{\rm S}(X)\ \ \alpha(h,x)\in\mbox{{\rm St}}_{\theta}\alpha(g,x)\ \mbox{\rm holds for}\ x\in\sigma,\ \mbox{\rm then}\\ \mbox{\rm there exists}\ \ g^{\prime}\in g\mbox{{\rm St}}_{\sigma}\ \ \mbox{\rm such that}\ \ h\in\mbox{{\rm St}}_{\theta}g^{\prime},\ \mbox{where, without loss of generality,}\ \sigma\subset\theta\in\Sigma_{X}.\\ \end{array}$
For any $g,h\in{\rm S}(X)$ such that $\alpha(h,x)\in\mbox{{\rm St}}_{\theta}\alpha(g,x)$, $x\in\sigma$, we consider, without loss of generality, the following.
Put $\sigma^{\prime}=\{x\in\sigma\ |\ \alpha(g,x)\in\theta\}\subset\sigma$. Then $\alpha(h,x)=\alpha(g,x)\in\theta$, $x\in\sigma^{\prime}$. If $\sigma^{\prime}\neq\sigma$, then $\alpha(g,x)\not\in\theta$, $\alpha(h,x)\not\in\theta$ for $x\in\sigma\setminus\sigma^{\prime}$.
If $\sigma^{\prime}=\sigma$, then $hg^{-1}\in\mbox{{\rm St}}_{\theta}$ and $h\in\mbox{{\rm St}}_{\theta}g$, $g\in g\mbox{{\rm St}}_{\sigma}$.
If $\sigma^{\prime}\neq\sigma$, then there exists $f\in{\rm S}(X)$ such that the points $x\in\theta,\alpha(g,y)$, where $y\in\sigma\setminus\sigma^{\prime}$, are mapped respectively to the points $x\in\theta,\alpha(h,y),y\in\sigma\setminus\sigma^{\prime}$ (for fixed orders on $\theta$ and $\sigma\setminus\sigma^{\prime}$). $f\in\mbox{{\rm St}}_{\theta}$ and $\alpha(h,x)=\alpha(fg,x),\ x\in\sigma$. Hence, $h^{-1}fg\in\mbox{{\rm St}}_{\sigma}$. Consequently $fg\in h\mbox{{\rm St}}_{\sigma}$ (equivalently $h\in(fg)\mbox{{\rm St}}_{\sigma}$) and there is $g^{\prime}\in g\mbox{{\rm St}}_{\sigma}$ such that $h\in\mbox{{\rm St}}_{\theta}g^{\prime}$. The inclusion $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\supset R_{\Sigma_{X}}$ and, hence, the equality $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{X}}$ are proved.
(c) At first, let us note that for any $f\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$ there are no points $x\neq y\in X$ such that $f(x)=f(y)\in X$. Indeed, suppose that there are $x\neq y\in X$ such that $f(x)=f(y)\in X$. One can take the nbd $W=\{h\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))\ |\ h(x)=h(y)=f(x)\}$ of $f$. Then $W\cap\tilde{\imath}({\rm S}(X))=\emptyset$ and $f\not\in\mbox{{\rm cl}}\tilde{\imath}({\rm S}(X))$.
Secondly, $f(\infty)=\infty$ for any $f\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$. Indeed, if $f(\infty)=x\in X$, then take the nbd $W=\{h\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))\ |\ h(\infty)=x\}$ of $f$ and $W\cap\tilde{\imath}({\rm S}(X))=\emptyset$ since $g(\infty)=\infty$ for any $g\in\tilde{\imath}({\rm S}(X))$.
It remains to check that $f\in b_{r}{\rm S}(X)$ if $f(\infty)=\infty$, $f$ is a bijection on $Y\subset X$ and $f(X\setminus Y)=\infty$, where $Y$ is an arbitrary subset of $X$. Any nbd of $f$ is of the form $\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$, $\sigma\in\Sigma_{X}$, where $Of(x)=f(x)$ if $f(x)\neq\infty$, and $Of(x)=\alpha X\setminus\sigma_{x}$, $\sigma_{x}\in\Sigma_{X}$, if $f(x)=\infty$. By finitness of $\sigma$ and $\sigma_{x}$, $x\in\sigma$ there is $h\in{\rm S}(X)$ such that $h(x)=f(x)$ if $x\in\sigma$ and $f(x)\in X$, and $h(x)\not\in\sigma_{x}$, if $x\in\sigma$ and $f(x)=\infty$. Evidently, $h\in\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$ and $f\in b_{r}{\rm S}(X)$.
(d) From (b) $b_{r}{\rm S}(X)=\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$ is a right topological semigroup, where $\tilde{\imath}$ is a topological isomorphism of ${\rm S}(X)$ into $(\alpha X)^{\alpha X}$, see § 2.2. If one shows that all elements of $\mbox{{\rm cl}}\tilde{\imath}({\rm S}(X))$ are continuous (as selfmaps of $\alpha X$), then multiplication on the left is continuous. To check the continuity of $f\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$ it is sufficient to check the continuity of $f$ at the point $\infty$.
Suppose that $f\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$ is not continuous at the point $\infty$. Then there exists a nbd $W=\alpha X\setminus\sigma$ of $\infty$ for some $\sigma\in\Sigma_{X}$ such that the set $\{x\in X\ |\ f(x)\in\sigma\}$ is infinite. Since $\sigma$ is finite, there exist $x\neq y\in X$ such that $f(x)=f(y)$. Therefore, $f\not\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$. The obtained contradiction shows that all $f\in\mbox{{\rm cl}}~\tilde{\imath}({\rm S}(X))$ are continuous and $b_{r}{\rm S}(X)$ is a semitopological semigroup. ∎
Remark 3.5.
(1) $b_{r}{\rm S}(X)$ is the set of maps $f:X\to\alpha X$ in the t.p.c. such that $f$ is a bijection on $Y\subset X$ and $f(X\setminus Y)=\infty$, where $Y$ is an arbitrary subset of $X$.
(2) In [18, § 12, Theorem 12.2] the description of $b_{r}{\rm S}(\mathbb{N})$ is given. It is also shown that $b_{r}{\rm S}(\mathbb{N})$ is a semitopological semigroup homeomorphic to the Cantor set.
Definition 3.6.
An action of a group $G$ on a set $X$ is strongly $n$-transitive, $n\geq 1$, if for any families of distinct $n$ points $x_{1},\ldots,x_{n}$ and $y_{1},\ldots,y_{n}$ there exists $g\in G$ such that $g(x_{k})=y_{k}$, $k=1,\ldots,n$.
An action $G\curvearrowright X$, which is strongly $n$-transitive for all $n\in\mathbb{N}$, is called ultratransitive.
Remark 3.7.
(1) A group $G$ which acts ultratransitively on $X$ is a dense subgroup of $({\rm S}(X),\tau_{\partial})$.
Indeed, let a set $O$ be open in $({\rm S}(X),\tau_{\partial})$ and $g\in O$. Then there are $\sigma\in\Sigma_{X}$ such that $g\mbox{{\rm St}}_{\sigma}\subset O$ and $g\mbox{{\rm St}}_{\sigma}=\{h\in G\ |\ h(x)=g(x),\ x\in\sigma\}$ is an open neighbourhood of $g$. Since $G$ acts ultratransitively on $X$, there is $h\in G$ such that $h(x)=g(x)$, $x\in\sigma$. Evidently, $h\in O$.
(2) The subgroup ${\rm S}_{<\omega}(X)$ of the group ${\rm S}(X)$ whose elements has finite supports acts ultratransitively on $X$. Hence, ${\rm S}_{<\omega}(X)$ is a dense subgroup of $({\rm S}(X),\tau_{\partial})$. The group whose elements have finite supports is a subgroup of any Houghton’s group (see, for example, [12]). Hence, Houghton’s groups act ultratransitively on the corresponding countable discrete spaces.
(3) Any group $G$ which acts ultratransitively on $X$ is oligomorphic and, hence, Roelcke precompact (see, [44], [42]). The Roelcke precompactness of $({\rm S}(X),\tau_{\partial})$ is proved in [16] (see also [37, Example 9.14]), the Roelcke precompactness of $({\rm S}_{<\omega}(X),\tau_{\partial})$ is proved in [4].
From (1) it follows that the Roelcke compactification of a group $G$ which acts ultratransitively on $X$ is isomorphic to the Roelcke compactification of $({\rm S}(X),\tau_{\partial})$.
(4) A space $X$ is ultrahomogeneous if the action of its homeomorphism group $\mbox{{\rm Hom}}(X)$ is ultratransitive. Locally compact metrizable CDH spaces whose complement to any finite subset is connected are ultrahomogeneous spaces. They include the spheres $S^{n-1}$ in Euclidean spaces $\mathbb{R}^{n}$, $n\geq 3$, the Hilbert cube $Q$ (see, for example, [2]).
The group of homeomorphisms of an ultrahomogeneous space with the permutation topology is Roelcke precompact, although the permutation topology is not necessarily admissible. The homeomorphism groups of the spheres $S^{n}$, $n\geq 2$, and the Hilbert cube $Q$ in the compact open topology (the smallest admissible group topology) are not Roelcke precompact [38]. Hence, in any admissible group topology, these groups are not Roelcke precompact.
Corollary 3.8.
Let $G$ acts ultratransitively on a discrete space $X$. Then for the action $(G,\tau_{\partial})\curvearrowright X$ items (a) and (b) of Theorem 3.4 hold and $b_{r}(G,\tau_{\partial})$ and $b_{r}{\rm S}(X)$ are topologically isomorphic semitopological semigroups.
Proof.
The same reasonings as in the proof of Theorem 3.4 work in the case of ultratransitive action to prove analogs of items (a) and (b).
$(G,\tau_{\partial})$ is a dense subgroup of $({\rm S}(X),\tau_{\partial})$ by item (1) of Remark 3.7 and is Roelcke precompact [37, Proposition 3.24]. Moreover, these Roelcke compacifications of $(G,\tau_{\partial})$ and $({\rm S}(X),\tau_{\partial})$ are uniformly isomorphic (use [37, Proposition 3.24] and [14, Corollary 8.3.11]) by the extension of the identity map on $(G,\tau_{\partial})$. Hence, $b_{r}(G,\tau_{\partial})$ is a semitopological semigroup topologically isomorphic to $b_{r}S(X)$. ∎
3.1.2. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultratransitive action on a chain
For an infinite chain (linearly ordered set) $X$ let $\mbox{{\rm Aut}}(X)$ be the group of automorphisms (order-preserving bijections) of $X$. A chain $X$ is called homogeneous if the action $\mbox{{\rm Aut}}(X)\curvearrowright X$ is transitive (for any $x,y\in X$ there exists $f\in\mbox{{\rm Aut}}(X)$ such that $f(x)=y$). A chain $X$ is called 2-homogeneous if for any pairs of points $x<y$ and $x^{\prime}<y^{\prime}$ there exists $g\in\mbox{{\rm Aut}}(X)$ such that $g(x)=x^{\prime}$, $g(y)=y^{\prime}$.
Remark 3.9.
(1) A 2-homogeneous chain is dense [27, Ch. 6, § 2] (and, hence, has no jumps).
(2) A 2-homogeneous chain is ultrahomogeneous [17, Lemma 1.10.1] (see also [35]), i.e. for any families of different $n$ points $x_{1}<\ldots<x_{n}$ and $y_{1}<\ldots<y_{n}$ there exists $g\in\mbox{{\rm Aut}}(X)$ such that $g(x_{k})=y_{k}$, $k=1,\ldots,n$, $n\in\mathbb{N}$.
(3) A chain without proper gaps is called continuous [27, Ch. 6, § 2]. If a chain $X$ is ultrahomogeneous and has a proper gap, then there is a gap on every nonempty interval of $X$ ($J\subset X$ is an interval if $x<y\in J\Longrightarrow z\in J\ \forall\ x<z<y$).
(4) On the group $\mbox{{\rm Aut}}(X)$, the permutation topology $\tau_{\partial}$ is the smallest admissible group topology for the action $\mbox{{\rm Aut}}(X)\curvearrowright X$, where $X$ is a discrete space. $(\mbox{{\rm Aut}}(X),\tau_{\partial})$ is a subgroup $({\rm S}(X),\tau_{\partial})$.
A chain $X$ in the discrete topology is a GO-space (generalized ordered space). There is the smallest LOTS
$X\otimes_{\ell}\{-1,0,1\}\ (\mbox{topology induced by the lexicographic order on}\ X\times\{-1,0,1\}),$
in which $X=X\times\{0\}$ is a dense subspace [34], and $X\otimes_{\ell}\{-1,0,1\}$ is naturally embedded in any other linearly ordered extension of $X$, in which $X$ is dense. Hence, any linearly ordered compactification of $X$ is a linearly ordered compactification of $X\otimes_{\ell}\{-1,0,1\}$. According to the description in [15], the smallest linearly ordered compactification [24] $b_{m}X$ of $X\otimes_{\ell}\{-1,0,1\}$ (and, hence, of $X$) is generated by replacing each gap (also improper) by a point with a natural continuation of the order (gaps in $X\otimes_{\ell}\{-1,0,1\}$ and $X$ are naturally identified). Therefore, $b_{m}X=X^{+}\cup X\cup X^{-}\cup\Gamma$, where $X^{+}=X\times\{1\}$, $X^{0}=X\times\{0\}$, $X^{-}=X\times\{-1\}$, $\Gamma$ is the set of gaps.
Remark 3.10.
There is the unique linearly ordered compactifications of $X$, when $X$ is a continuous chain. It is obtained by addition points $\sup$ and $\inf$ to $X\otimes_{\ell}\{-1,0,1\}$.
Theorem 3.11.
Let $X$ be an ultrahomogeneous chain and a discrete space, $G=(\mbox{{\rm Aut}}(X),\tau_{\partial})$, $\mathcal{U}_{X}$ is the maximal equiuniformity on $X$ for the action $\alpha:G\times X\to X$. Then
  • (a)
    $\mathcal{U}_{X}\in\mathbb{BU}(X)$ and $\tilde{X}=b_{m}X$, where $(\tilde{X},\tilde{\mathcal{U}}_{X})$ is the completion of $(X,\mathcal{U}_{X})$,
  • (b)
    $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=L\wedge R=R_{\Sigma_{X}}\subset R_{\Sigma_{b_{m}X}}$ (and, hence, $G$ is Roelcke precompact),
  • (c)
    $L\wedge R\subset\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{b_{m}X}}$ (and, hence, $\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)=\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{X})$ and $L\wedge R\not\in\mathbb{E}(G)$ if $R_{\Sigma_{X}}\neq R_{\Sigma_{b_{m}X}}$),
  • (d)
    the proper Ellis semigroup compactification $bG$ (the completion of $(G,\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)=R_{\Sigma_{b_{m}X}})$) is the set of selfmaps $f$ of $b_{m}X$ in the t.p.c. such that
    • (i)
      $f$ is monotone (if $x<y$, then $f(x)\leq f(y)$),
    • (ii)
      $f(x)\not\in X$ if $x\in b_{m}X\setminus X$,
    • (iii)
      if $f(x)=f(y)$, $x\neq y$, then $f(x)\in b_{m}X\setminus X$,
    • (iv)
      either $f((x,-1))=f((x,0))=f((x,1))\not\in X$, or $f((x,-1))=(y,-1)$, $f((x,0))=(y,0)$, $f((x,1))=(y,1)$ and
    • (v)
      $f(\inf)=\inf$, $f(\sup)=\sup$,
  • (e)
    the Roelcke compactification $b_{r}G$ is the set of maps $f$ of $X$ to $b_{m}X$ in the t.p.c. such that (i’) $f$ is monotone, (ii’) if $f(x)=f(y)$, $x\neq y$, then $f(x)\in b_{m}X\setminus X$.
If $X$ is continuous, then
  • (b’)
    $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=L\wedge R=R_{\Sigma_{X}}=R_{\Sigma_{b_{m}X}}$,
  • (c’)
    $b_{r}G=bG$ is a proper enveloping Ellis semigroup compactification.
Proof.
(a) Firstly, the base of the maximal equiuniformity $\mathcal{U}_{X}$ for the action $(G,\tau_{\partial})\curvearrowright X$ is formed by the finite covers (of disjoint sets)
$u_{\theta}=\{\mbox{{\rm St}}_{\theta}x\ |\ x\in X\}=(\leftarrow,x_{1})\cup\{x_{1}\}\cup(x_{1},x_{2})\cup\{x_{2}\}\cup\ldots\cup(x_{n-1},x_{n})\cup\{x_{n}\}\cup(x_{n},\to),\ \theta\in\Sigma_{X},$
($\theta=\{x_{1},\ldots,x_{n}\}$, $x_{1}<\ldots<x_{n}\in X$, $u_{\theta^{\prime}}\succ u_{\theta}\ \mbox{if}\ \theta\subset\theta^{\prime}$). Hence, $\mathcal{U}_{X}\in\mathbb{BU}(X)$.
Secondly, the smallest linearly ordered compactification $b_{m}X$ of $X$ is zero-dimensional. In any open cover of a zero-dimensional compactum $b_{m}X$ it is possible to refine a finite cover of pairwise distinct clopen intervals no none is a subset of the other. By introducing the order on the intervals ($[a_{1},b_{1}]<[a_{2},b_{2}]$ if $a_{1}<a_{2}$), we obtain a disjoint cover of the clopen intervals $J_{1}=(\leftarrow,y_{1}),\ldots,J_{k}=(y_{k},\to)$, subtracting sequentially from the $i$ interval the union of the previous ones. Let us correct the latter cover by removing endpoints of the form $(x,0)$ from the intervals (if any), and adding them as single-point clopen intervals to the corrected interval system to obtain an open cover $\omega$ of the space $b_{m}X$. Then $\omega\wedge X$ (the trace of $\omega$ on $X$) obviously coincides with the cover from $\mathcal{U}_{X}$. Since any cover of $\mathcal{U}_{X}$ is extended to the cover of $b_{m}X$ ($u_{\theta}$ is a trace of the cover $\omega_{\theta}=[\inf,(x_{1},-1)]\cup\{(x_{1},0)\}\cup[(x_{1},1),(x_{2},-1)]\cup\{(x_{2},0)\}\cup\ldots\cup[(x_{n-1},1),(x_{n},-1)]\cup\{(x_{n},0)\}\cup[(x_{n},1),\sup]$), then $\tilde{X}=b_{m}X$ and the action $\alpha:G\times X\to X$ is extended to the action $\tilde{\alpha}:G\times b_{m}X\to b_{m}X$ (for $g\in G$
  • $\tilde{\alpha}(g,(x,j))=(\alpha(g,x),j)$, $x\in X$, $j=-1,0,1$;
  • $\tilde{\alpha}(g,(A,B))=(\alpha(g,A),\alpha(g,B))$, $(A,B)\in\Gamma$).
It is easy to check that the action $\tilde{\alpha}$ is correctly defined, it is continuous, its restriction to $X=X^{0}$ coincides with $\alpha$ and $X^{+}$, $X^{0}$, $X^{-}$, $\Gamma$ are invariant subsets. If $\Omega_{\theta}$ is an entourage correspondent to the cover $\omega_{\theta}$, then $\Omega_{\theta}\circ\Omega_{\theta}=\Omega_{\theta}$.
The equality $L\wedge R=R_{\Sigma_{X}}$ and the inclusion $R_{\Sigma_{X}}\subset R_{\Sigma_{b_{m}X}}$ in (b) follow from Lemma 3.3 and Theorem 3.2 respectively.
Proof of the equality $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{b_{m}X}}$ in (c). By Proposition 2.8 $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\subset R_{\Sigma_{b_{m}X}}$.
Since $\mbox{{\rm St}}_{(x,j)}$, $j=-1,0,1$, for the action $\tilde{\alpha}$ coincides with $\mbox{{\rm St}}_{x}$ for the action $\alpha$ and $\mbox{{\rm St}}_{\inf}=\mbox{{\rm St}}_{\sup}=G$, one can take arbitrary $\sigma\in\Sigma_{b_{m}X}$, where $\sigma=\sigma_{X}\cup\sigma_{\Gamma}$, $\sigma_{X}\in\Sigma_{X^{0}}(=\Sigma_{X})$, $\sigma_{\Gamma}\in\Sigma_{\Gamma}\setminus\{\inf,\sup\}$.
[…]
$\begin{array}[]{c}\mbox{\rm if for}\ g,h\in G\ \ (\tilde{\alpha}(h,x),\tilde{\alpha}(g,x))\in\Omega_{\theta}\ \mbox{\rm holds for}\ x\in\sigma=\sigma_{X}\cup\sigma_{\Gamma},\\ \mbox{\rm then there exists}\ \ g^{\prime}\in g\mbox{{\rm St}}_{\sigma}\ \ \mbox{\rm such that}\ (\tilde{\alpha}(h,x),\tilde{\alpha}(g^{\prime},x))\in\Omega_{\theta}\ \mbox{\rm holds for}\ x\in b_{m}X,\\ \mbox{where, without loss of generality,}\ \sigma_{X}\subset\theta\in\Sigma_{X}.\\ \end{array}$
Let $\sigma=\{y_{1},\ldots,y_{n}\}$ and $y_{1}<y_{2}<\ldots<y_{n}$. The construction of the automorphism $g^{\prime}$ is carried out on the intervals $[\inf,y_{1}],[y_{1},y_{2}],\ldots,[y_{n-1},y_{n}],[y_{n},\sup]$.
(I) The map $g^{\prime}$ on the interval $[\inf,y_{1}]$. If $(\inf,\tilde{\alpha}(h,y_{1}))\in\Omega_{\theta}$ or $\tilde{\alpha}(h,y_{1})\in\theta$ (and hence, $\tilde{\alpha}(g,y_{1})=\tilde{\alpha}(h,y_{1})$, $y_{1}=(x_{1},0)$, $\tilde{\alpha}(g,(x_{1},-1))=\tilde{\alpha}(h,(x_{1},-1))$) and $(\inf,\tilde{\alpha}(h,(x_{1},-1)))\in\Omega_{\theta}$, then
$\tilde{\alpha}(g^{\prime},y)=\tilde{\alpha}(g,y)\ \mbox{for}\ y\in[\inf,y_{1}].$
Otherwise, in the case $\tilde{\alpha}(g,y_{1})\leq\tilde{\alpha}(h,y_{1})$ if $\tilde{\alpha}(g,y_{1})\not\in\theta$, then take $a_{1}<\tilde{\alpha}(g,y_{1})$ such that $a_{1}\in X^{0}$, $(a_{1},\tilde{\alpha}(g,y_{1}))\in\Omega_{\theta}$ (and hence, $(a_{1},\tilde{\alpha}(h,y_{1}))\in\Omega_{\theta}$). If $\tilde{\alpha}(g,y_{1})\in\theta$, then $y_{1}=(x_{1},0)$, $\tilde{\alpha}(g,y_{1})=\tilde{\alpha}(h,y_{1})$ and $\tilde{\alpha}(g,(x_{1},-1))=\tilde{\alpha}(h,(x_{1},-1))$. Take $a_{1}<\tilde{\alpha}(g,(x_{1},-1))$ such that $a_{1}\in X^{0}$, $(a_{1},\tilde{\alpha}(g,(x_{1},-1)))\in\Omega_{\theta}$ (and hence, $(a_{1},\tilde{\alpha}(h,(x_{1},-1)))\in\Omega_{\theta}$). Put $t=y_{1}$ if $\tilde{\alpha}(g,y_{1})\not\in\theta$ and $t=(x_{1},-1)$ if $\tilde{\alpha}(g,y_{1})\in\theta$.
We assume that $y_{1}^{-}=\alpha(h^{-1},a_{1})$ and take $y_{1}^{-}<z_{1}^{-}<t$, $z_{1}^{-}\in X^{0}$, $a_{1}<\alpha(g,z_{1}^{-})$ (using continuity of $g$ and the inequality $a_{1}<\tilde{\alpha}(g,t)$). Since $X$ is ultrahomogeneous, there exists $\varphi_{1}\in G$ such that $\alpha(\varphi_{1},y_{1}^{-})=a_{1}=\alpha(h,y_{1}^{-})<\alpha(\varphi_{1},z_{1}^{-})=\alpha(g,z_{1}^{-})$.
In the case $\tilde{\alpha}(g,y_{1})>\tilde{\alpha}(h,y_{1})$ take $a_{1}<\tilde{\alpha}(h,y_{1})$ such that $a_{1}\in X^{0}$, $(a_{1},\tilde{\alpha}(h,y_{1}))\in\Omega_{\theta}$ (and hence, $(a_{1},\tilde{\alpha}(g,y_{1}))\in\Omega_{\theta}$). We assume that $y_{1}^{-}=\alpha(h^{-1},a_{1})$ and take $y_{1}^{-}<z_{1}^{-}<y_{1}$, $z_{1}^{-}\in X^{0}$, $a_{1}<\alpha(g,z_{1}^{-})$ (using continuity of $g$ and the inequality $a_{1}<\tilde{\alpha}(g,y_{1})$). Since $X$ is ultrahomogeneous, there exists $\varphi_{1}\in G$ such that $\alpha(\varphi_{1},y_{1}^{-})=a_{1}=\alpha(h,y_{1}^{-})<\alpha(\varphi_{1},z_{1}^{-})=\alpha(g,z_{1}^{-})$. Put
$\tilde{\alpha}(g^{\prime},y)=\left\{\begin{array}[]{lcl}\tilde{\alpha}(h,y)&\mbox{if}&y\in[\inf,y_{1}^{-}],\\ \tilde{\alpha}(\varphi_{1},y)&\mbox{if}&y\in[y_{1}^{-},z_{1}^{-}],\\ \tilde{\alpha}(g,y)&\mbox{if}&y\in[z_{1}^{-},y_{1}].\\ \end{array}\right.$
On the interval $[y_{n},\sup]$ the construction is similar.
(II) The map $g^{\prime}$ on on the interval $[y_{k},y_{k+1}]$, $k=1,\dots,n-1$. If
$(\tilde{\alpha}(h,y_{k}),\tilde{\alpha}(h,y_{k+1}))\in\Omega_{\theta},\ \mbox{or}$
$\tilde{\alpha}(h,y_{k})\in\theta$, $\tilde{\alpha}(h,y_{k+1})\in\theta$ (and hence, $\tilde{\alpha}(g,y_{k})\in\theta$, $\tilde{\alpha}(g,y_{k+1})\in\theta$, $y_{k}=(x_{k},0)$, $y_{k+1}=(x_{k+1},0)$, $\tilde{\alpha}(h,(x_{k},1))=\tilde{\alpha}(g,(x_{k},1))$, $\tilde{\alpha}(h,(x_{k+1},-1))=\tilde{\alpha}(g,(x_{k+1},-1))$) and $(\tilde{\alpha}(h,(x_{k},1)),\tilde{\alpha}(h,(x_{k+1},-1)))\in\Omega_{\theta}$, or
$\tilde{\alpha}(h,y_{k})\in\theta$, $\tilde{\alpha}(h,y_{k+1})\not\in\theta$ and $(\tilde{\alpha}(h,(x_{k},1)),\tilde{\alpha}(h,y_{k+1}))\in\Omega_{\theta}$, or
$\tilde{\alpha}(h,y_{k})\not\in\theta$, $\tilde{\alpha}(h,y_{k+1})\in\theta$ and $(\tilde{\alpha}(h,y_{k}),\tilde{\alpha}(h,(x_{k+1},-1)))\in\Omega_{\theta}$, then
$\tilde{\alpha}(g^{\prime},y)=\tilde{\alpha}(g,y)\ \mbox{for}\ y\in[y_{k},y_{k+1}].$
Otherwise, let us examine the case $\tilde{\alpha}(h,y_{k})\geq\tilde{\alpha}(g,y_{k})$, $\tilde{\alpha}(h,y_{k+1})\geq\tilde{\alpha}(g,y_{k+1})$.
(a) If $\tilde{\alpha}(h,y_{k}),\tilde{\alpha}(h,y_{k+1})\not\in\theta$ (and, hence, $\tilde{\alpha}(g,y_{k}),\tilde{\alpha}(g,y_{k+1})\not\in\theta$), then take $b_{k}>\tilde{\alpha}(h,y_{k})$, $b_{k}\in X^{0}$, $(b_{k},\tilde{\alpha}(h,y_{k}))\in\Omega_{\theta}$ and $a_{k+1}<\tilde{\alpha}(g,y_{k+1})$, $a_{k+1}\in X^{0}$, $(a_{k+1},\tilde{\alpha}(g,y_{k+1}))\in\Omega_{\theta}$. $b_{k}<a_{k+1}$. Put $t=y_{k}$ and $\tau=y_{k+1}$.
(b) If $\tilde{\alpha}(h,y_{k})$, $\tilde{\alpha}(h,y_{k+1})\in\theta$, then take $b_{k}>\tilde{\alpha}(h,(x_{k},1))$, $b_{k}\in X^{0}$, $(b_{k},\tilde{\alpha}(h,(x_{k},1)))\in\Omega_{\theta}$ and $a_{k+1}<\tilde{\alpha}(g,(x_{k+1},-1))$, $a_{k+1}\in X^{0}$, $(a_{k+1},\tilde{\alpha}(g,(x_{k+1},-1)))\in\Omega_{\theta}$. $b_{k}<a_{k+1}$. Put $t=(x_{k},1)$ and $\tau=(x_{k+1},-1)$.
(c) If $\tilde{\alpha}(h,y_{k})\in\theta$, $\tilde{\alpha}(h,y_{k+1})\not\in\theta$, then take $b_{k}>\tilde{\alpha}(h,(x_{k},1))$, $b_{k}\in X^{0}$, $(b_{k},\tilde{\alpha}(h,(x_{k},1)))\in\Omega_{\theta}$ and $a_{k+1}<\tilde{\alpha}(g,y_{k+1})$, $a_{k+1}\in X^{0}$, $(a_{k+1},\tilde{\alpha}(g,y_{k+1}))\in\Omega_{\theta}$. $b_{k}<a_{k+1}$. Put $t=(x_{k},1)$ and $\tau=y_{k+1}$.
(d) If $\tilde{\alpha}(h,y_{k})\not\in\theta$, $\tilde{\alpha}(h,y_{k+1})\in\theta$, take $b_{k}>\tilde{\alpha}(h,y_{k})$, $b_{k}\in X^{0}$, $(b_{k},\tilde{\alpha}(h,y_{k}))\in\Omega_{\theta}$ and $a_{k+1}<\tilde{\alpha}(g,(x_{k+1},-1))$, $a_{k+1}\in X^{0}$, $(a_{k+1},\tilde{\alpha}(g,(x_{k+1},-1)))\in\Omega_{\theta}$. $b_{k}<a_{k+1}$. Put $t=y_{k}$ and $\tau=(x_{k+1},-1)$.
We assume that $y_{k}^{+}=\alpha(h^{-1},b_{k})$, $y_{k+1}^{-}=\alpha(h^{-1},a_{k+1})$. From the continuity of $g$ and the inequality $\tilde{\alpha}(g,t)<b_{k}$ there exists $t<z^{+}_{k}<y_{k}^{+}$, $z^{+}_{k}\in X^{0}$, such that $\alpha(g,z_{k}^{+})<b_{k}$, and from the inequality $a_{k+1}<\tilde{\alpha}(g,\tau)$ there exists $y_{k+1}^{-}<z^{-}_{k+1}<\tau$, $z^{-}_{k+1}\in X^{0}$, such that $a_{k+1}<\alpha(g,z_{k+1}^{-})$. Since $X$ is ultrahomogeneous, there exist $\varphi_{k}^{-}\in G$ such that $\alpha(\varphi_{k}^{-},z^{+}_{k})=\alpha(g,z_{k}^{+})$, $\alpha(\varphi_{k}^{-},y_{k}^{+})=b_{k}=\alpha(h,y_{k}^{+})$ and $\varphi_{k}^{+}\in G$ such that $\alpha(\varphi_{k}^{+},y_{k+1}^{-})=a_{k+1}=\alpha(h,y_{k+1}^{-})<\alpha(\varphi_{k}^{+},z^{-}_{k+1})=\alpha(g,z_{k+1}^{-})$. Put
$g^{\prime}(x)=\left\{\begin{array}[]{lcl}\tilde{\alpha}(g,y)&\mbox{if}&y\in[y_{k},z_{k}^{+}],\\ \tilde{\alpha}(\varphi_{k}^{-},y)&\mbox{if}&y\in[z_{k}^{+},y_{k}^{+}],\\ \tilde{\alpha}(h,y)&\mbox{if}&y\in[y_{k}^{+},y_{k+1}^{-}],\\ \tilde{\alpha}(\varphi_{k}^{+},y)&\mbox{if}&x\in[y_{k+1}^{-},z_{k+1}^{-}],\\ \tilde{\alpha}(g,y)&\mbox{if}&y\in[z_{k+1}^{-},y_{k+1}].\\ \end{array}\right.$
In the cases $\tilde{\alpha}(h,y_{k})\leq\tilde{\alpha}(g,y_{k})$, $\tilde{\alpha}(h,y_{k+1})\leq\tilde{\alpha}(g,y_{k+1})$, $\tilde{\alpha}(h,y_{k})\geq\tilde{\alpha}(g,y_{k})$, $\tilde{\alpha}(h,y_{k+1})\leq\tilde{\alpha}(g,y_{k+1})$ and $\tilde{\alpha}(h,y_{k})\leq\tilde{\alpha}(g,y_{k})$, $\tilde{\alpha}(h,y_{k+1})\geq\tilde{\alpha}(g,y_{k+1})$ considerations are similar.
[…]
From item (4) of Theorem 3.2 it follows that $\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)=\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{X})$.
If $R_{\Sigma_{X}}\neq R_{\Sigma_{b_{m}X}}$, then $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})\neq\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$, $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=L\wedge R\neq\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)=\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{X})$. Hence, by item (e) of Theorem 2.22 $L\wedge R\not\in\mathbb{E}(G)$ and (c) is proved.
(d) Fulfillment of conditions (i)–(v) for $f\in\mbox{{\rm cl}}~G$.
(i) Let $x<y$ and $f(x)>f(y)$. Take disjoint nbds $Of(x)$, $Of(y)$ of $f(x)$ and $f(y)$ respectively, which are intervals. Then $O_{f}=[x,Of(x)]\cap[y,Of(y)]$ is a nbd of $f$, but $G\cap O_{f}=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~G$.
(ii) Let $x\in b_{m}X\setminus X$ and $f(x)\in X$. Then $O_{f}=[x,\{f(x)\}]$ is a nbd of $f$, but $G\cap O_{f}=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~G$.
(iii) Let $f(x)=f(y)$, $x\neq y$, and $f(x)\in X$. Then $O_{f}=[x,\{f(x)\}]\cap[y,\{f(y)\}]$ is a nbd of $f$, but $G\cap O_{f}=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~G$.
(iv) If $f((x,-1))=f((x,0))<f((x,1))$, then $f((x,0))$, $f((x,1))\in b_{m}X\setminus X$ by (ii) and then there exists $f((x,0))<z<f((x,1))$. For the disjoint nbds $U=[\inf,z)$ and $V=(z,\sup]$ of $f((x,0))$ and $f((x,1))$ respectively, $O_{f}=[(x,0),U]\cap[(x,1),V]$ is a nbd of $f$, but $G\cap O_{f}=\emptyset$.
This argument works in the case $f((x,-1))<f((x,0))=f((x,1))$. Therefore, either $f((x,-1))=f((x,0))=f((x,1))\not\in X$ or $f((x,-1))<f((x,0))<f((x,1))$ and there is no $z$ (by the above argument) such that $f((x,-1))<z<f((x,0))$ or $f((x,0)<z<f((x,1))$. Hence, $f((x,-1))=(y,-1)$, $f((x,0))=(y,0)$, $f((x,1))=(y,1)$.
(v) Let $f(\inf)=x\neq\inf$. Take nbd $Ox$ of $x$, $\inf\not\in Ox$. Then $O_{f}=[\inf,Ox]$ is a nbd of $f$, but $G\cap O_{f}=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~G$. The same argument works for $\sup$.
[…]
In order to construct automorphism $g\in O_{f}$ we sequentially determine two finite increasing sequences of points $t_{k}\leq t^{\prime}_{k}<t_{k+1}$ and $\tau_{k}\leq\tau^{\prime}_{k}<\tau_{k+1}$ in $X$, $k=1,\ldots,n$.
If $Of(x_{k})=\{(y_{k},0)\}$, then $x_{k}=(t_{k},0)$ by (ii), $t^{\prime}_{k}=t_{k}$, and $\tau_{k}=\tau^{\prime}_{k}=y_{k}$.
Let $Of(x_{k})=[(y_{k},1),(z_{k},-1)]$. If
  • ($\alpha$)
    $x_{k}=(r_{k},-1)$ and $f(x_{k})\neq(z_{k},-1)$, or $f((r_{k},0))=(z_{k},-1)$ ($f((r_{k},-1))=f((r_{k},1))=(z_{k},-1)$ by (iv)),
  • ($\beta$)
    $x_{k}=(r_{k},0)$,
  • ($\gamma$)
    $x_{k}=(r_{k},1)$ and $f(x_{k})\neq(y_{k},1)$, or $f((r_{k},0))=(y_{k},1)$ ($f((r_{k},-1))=f((r_{k},1))=(z_{k},1)$ by (iv)),
then take any $(\tau_{k},0)\in[(y_{k},1),(z_{k},-1)]$, $\tau^{\prime}_{k-1}<\tau_{k}$ (if $k-1=0$ , then $\inf<(\tau_{k},0)$) and put $t_{k}=t^{\prime}_{k}=r_{k}$, $\tau_{k}=\tau^{\prime}_{k}$.
If $x_{k}=(r_{k},-1)$, $f(x_{k})=(z_{k},-1)$ and $f((r_{k},0))=(z_{k},0)$ (by (iv)), then put $t_{k}=t^{\prime}_{k}=r_{k}$, $\tau_{k}=\tau^{\prime}_{k}=z_{k}$.
If $x_{k}=(r_{k},1)$, $f(x_{k})=(y_{k},1)$ and $f((r_{k},0))=(y_{k},0)$ (by (iv)), then put $t_{k}=t^{\prime}_{k}=r_{k}$, $\tau_{k}=\tau^{\prime}_{k}=y_{k}$.
If $x_{k}\in\Gamma$, then take any $(\tau_{k},0),\ (\tau^{\prime}_{k},0)\in[(y_{k},1),(z_{k},-1)]$, $(\tau^{\prime}_{k-1},0)<(\tau_{k},0)$ and any $t_{k},\ t^{\prime}_{k}\in X$ such that $(t^{\prime}_{k-1},0)<(t_{k},0)<x_{k}<(t^{\prime}_{k},0)<x_{k+1}$ (if $k-1=0$ , then $\inf<(t_{k},0)$; if $k=n$, then $(t^{\prime}_{k},0)<\sup$).
Since $X$ is ultrahomogeneous there exists $g\in G$ that maps ordered sequence of $t_{k},\ t^{\prime}_{k}$ onto the ordered sequence of $\tau_{k},\ \tau^{\prime}_{k}$. Evidently, $g\in O_{f}$ and $f\in\mbox{{\rm cl}}~G$.
(e) Since $L\wedge R=R_{\Sigma_{X}}$, the Roelcke compactification $b_{r}G$ of $G$ is the closure of $\imath(G)$ in $(b_{m}X)^{X}$. The restriction of the projection $\mbox{{\rm pr}}:(b_{m}X)^{b_{m}X}\to(b_{m}X)^{X}$ to $\mbox{{\rm cl}}~\tilde{\imath}(G)$ is a map of the proper Ellis semigroup compactification $bG$ onto $b_{r}G$. This gives the description of $b_{r}G$ as the set of maps $f$ of $X$ to $b_{m}X$ in the t.p.c. such that (a) $f$ is monotone, (b) if $f(x)=f(y)$, then $f(x)\in b_{m}X\setminus X$.
If $X$ is continuous, then (b’) follows from the equalities $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=\mathfrak{E}_{\alpha}(\mathcal{U}_{X})$ and $R_{\Sigma_{X}}=R_{\Sigma_{b_{m}X}}$.
$R_{\Sigma_{X}}=R_{\Sigma_{b_{m}X}}$, because $\mbox{{\rm St}}_{\inf}={\rm St}_{\sup}=G$ and $\mbox{{\rm St}}_{(x,-1)}=\mbox{{\rm St}}_{(x,1)}=\mbox{{\rm St}}_{(x,0)}$ for $x\in X$.
Since $\tilde{\alpha}(g,\inf)=\inf$, $\tilde{\alpha}(g,\sup)=\sup$, $g\in G$, in order to prove $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=\mathfrak{E}_{\alpha}(\mathcal{U}_{X})$ it is enough to check the fulfillment of the condition ()($\star$)( ⋆ ) of Proposition 2.14. Let us show that for any $\Omega\in\tilde{\mathcal{U}}_{X}$ and any point $x\in X$ the following hold
$(\tilde{\alpha}(g,(x,-1)),\tilde{\alpha}(f,(x,-1))\in\Omega\ \mbox{iff}\ (\tilde{\alpha}(g,(x,1)),\tilde{\alpha}(f,(x,1))\in\Omega\ \mbox{iff}$
$(\tilde{\alpha}(g,(x,0)),\tilde{\alpha}(f,(x,0))\in\Omega.$
Let, for example, $(x,1)\in b_{m}X$ and an arbitrary cover $\omega$ of $\{(x,1)\}\times b_{m}X$ from $\tilde{\mathcal{U}}_{X}$ be of the form
$\omega=[\inf,(x_{1},-1)]\bigcup\{(x_{1},0)\}\bigcup[(x_{1},1),(x_{2},-1)]\bigcup\{(x_{2},0)\}\bigcup\ldots$
$\ldots\bigcup[(x_{n-1},1),(x_{n},-1)]\bigcup\{(x_{n},0)\}\bigcup[(x_{n},1),\sup],$
where $x_{1}<x_{2}<\ldots<x_{n-1}<x_{n}$. It consists of disjoint sets. Let us note that if $(y,0)\in[\inf,(x_{1},-1)]$ or $[(x_{k},1),(x_{k+1},-1)]$, $k=1,\ldots,n-1$, or $[(x_{n},1),\sup]$, then $(y,-1),(y,1)\in[\inf,(x_{1},-1)]$ or $[(x_{k},1),(x_{k+1},-1)]$, $k=1,\ldots,n-1$, or $[(x_{n},1),\sup]$ respectively.
Take $(x,0)\in b_{m}X$ and the same cover $\omega$ on $\{(x,0)\}\times b_{m}X$.
If $f(x,0)=g(x,0)$, then $f(x,1)=g(x,1)$. In particular, if $f(x,0)=g(x,0)=(x_{k},0)$, then $f(x,1)=g(x,1)=(x_{k},1)\in[(x_{k},1),(x_{k+1},-1)]$, $k=1,\ldots,n$.
[…]
In the case $(x,-1)\in b_{m}X$ the same reasoning can be used and (b’) is proved.
(c’) follows from (b’). ∎
Remark 3.12.
$bG$ is not a semitopological semigroup.
Indeed, take the map $f\in bG$, $f(\inf)=\inf$, $f(x)=\sup$ otherwise. Any nbd of the map $g\in bG$, $g(\sup)=\sup$, $g(x)=\inf$ otherwise, is of the form $O_{x}^{y}=\{h\in bG\ |\ h((x,0))<(y,0)\}$, $x,y\in X$. Since $f\circ g=g$ in any nbd of $g$ of the form of $O_{x}^{y}$ there is $h\in bG$ such that $h(\inf)=\inf$, $h(x)=(y,-1)$, $\inf<x<\sup$, $h(\sup)=\sup$. Hence, $f\circ h(\inf)=\inf$, $f\circ h(x)=\sup$ and can’t belong to the arbitrary nbd of $g$ the form $O_{x}^{y}$. Thus, multiplication on the left in $bG$ is not continuous.
Corollary 3.13.
Let $X$ be an ultrahomogeneous chain and a discrete space, $G=(\mbox{{\rm Aut}}(X),\tau_{\partial})$. Then
  • (1)
    there is no $\mathcal{U}\in\mathbb{E}(G)$ such that $L\wedge R\subsetneq\mathcal{U}\subsetneq\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{X})$,
  • (2)
    $L\wedge R\in\mathbb{E}(G)$ iff $X$ is continuous.
Proof.
(1) follows from the equality $\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)=\widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{X})$ (item (c) of Theorem 3.11) and items (d), (f) of Theorem 2.22.
Sufficiency in (2) is proved in item (b’) of Theorem 3.11. Necessity. Suppose that $X$ is not continuous. It is enough to show that $R_{\Sigma_{X}}\neq R_{\Sigma_{b_{m}X}}$ by item (c) of Theorem 3.11.
Take $x\in X$ and let $\omega=[\inf,(x,-1)]\cup\{(x,0)\}\cup[(x,1),\sup]\in\tilde{\mathcal{U}}_{X}$, $\xi$ is a proper gap in $X$, $O\in N_{G}(e)$ is such that $\{Ox\ |\ x\in b_{m}X\}\succ\omega$.
For any $\sigma\in\Sigma_{X}$ let $(x_{1},0)<\xi<(x_{2},0)$, $x_{1},x_{2}\in\sigma$ and $\sigma\cap((x_{1},0),(x_{2},0))=\emptyset$. Take $y_{1},y_{2},y_{1}^{\prime},y_{2}^{\prime}\in X$ such that $(x_{1},0)<(y_{1},0)<(y_{1}^{\prime},0)<\xi<(y_{2},0)<(y_{2}^{\prime},0)<(x_{2},0)$ and $g\in G$ such that $g((y_{2},0))<(x,0)$, $g((y_{2}^{\prime},0))>(x,0)$. Let $h\in G$ sends points $x_{1}<y_{2}<y_{2}^{\prime}<x_{2}$ to $x_{1}<y_{1}<y_{1}^{\prime}<x_{2}$ and identity on the points from $\sigma$. Then $h\in\mbox{{\rm St}}_{\sigma}$ and, hence, for any $V\in N_{G}(e)$ $gh\in Vg\mbox{{\rm St}}_{\sigma}$. Further, $g(\xi)<(x,0)$, $gh(\xi)>(x,0)$. Therefore, $gh\not\in Og\mbox{{\rm St}}_{\xi}$.
It is shown that for any $\sigma\in\Sigma_{X}$ and $V\in N_{G}(e)$ the cover $\{Vg\mbox{{\rm St}}_{\sigma}\ |\ g\in G\}$ is not refined in the cover $\{Og\mbox{{\rm St}}_{\xi}\ |\ g\in G\}$. Hence, $L\wedge R=R_{\Sigma_{X}}\neq R_{\Sigma_{b_{m}X}}=\widetilde{{\mathfrak{E}}}_{\star}(L\wedge R)$. ∎
Remark 3.14.
Corollary 3.13 shows that an oligomorphic action (see [44] and [42]) doesn’t guarantee that the Roelcke compactification of an acting group is a proper Ellis semigroup compactification.
Definition 3.15.
An action of a group $G\subset\mbox{{\rm Aut}}X$ on a chain $X$ is strongly $n$-transitive, $n\geq 1$, if for any families of $n$ points $x_{1}<\ldots<x_{n}$ and $y_{1}<\ldots<y_{n}$ there exists $g\in G$ such that $g(x_{k})=y_{k}$, $k=1,\ldots,n$.
An action $G\curvearrowright X$, which is strongly $n$-transitive for all $n\in\mathbb{N}$, is called ultratransitive.
Remark 3.16.
If an action of $G$ on a chain $X$ is ultratransitive, then $(G,\tau_{\partial})$ is a dense subgroup $(\mbox{{\rm Aut}}(X),\tau_{\partial})$ and $X$ is a coset space of $(G,\tau_{\partial})$.
Indeed, for any $g\in\mbox{{\rm Aut}}(X)$ and a nbd $\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$ of $g$, ultratransitivity of the action $G\curvearrowright X$ yields that there exists $f\in G\cap\{h\in\mbox{{\rm Aut}}(X)\ |\ h(x_{i})=g(x_{i}),\ i=1,\ldots,n\}$.
Corollary 3.17.
Let an action $\alpha_{G}:(G,\tau_{\partial})\times X\to X$ on a chain $X$ (discrete space) be ultratransitive, $\mathcal{U}_{X}^{G}$ is the maximal equiuniformity on $X$. Then
  • (a)
    $\mathcal{U}_{X}=\mathcal{U}_{X}^{G}$, where $\mathcal{U}_{X}$ is the maximal equiuniformity on $X$ for the action $\alpha:(\mbox{{\rm Aut}}(X),\tau_{\partial})\times X\to X$,
  • (b)
    $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})|_{G}=\mathfrak{E}_{\alpha_{G}}(\mathcal{U}_{X}^{G})=(L\wedge R)_{G}$ and $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})|_{G}=\widetilde{\mathfrak{E}}_{\alpha_{G}}(\mathcal{U}_{X}^{G})$, where $(L\wedge R)_{G}$ is the Roelcke uniformity on $G$,
  • (c)
    $\widetilde{\mathfrak{E}}_{\alpha_{G}}(\mathcal{U}_{X}^{G})=\widetilde{{\mathfrak{E}}}_{\star}((L\wedge R)_{G})$,
  • (d)
    $bG$ and $b\,\mbox{{\rm Aut}}(X)$ are topologically isomorphic proper enveloping Ellis semigroup compactifications.
If $X$ is continuous, then the Roelcke compactifications $b_{r}G$ and $b_{r}\,\mbox{{\rm Aut}}(X)$ are topologically isomorphic proper enveloping Ellis semigroup compactifications. If the Roelcke compactifications $b_{r}G$ is a proper enveloping Ellis semigroup compactification, then $X$ is continuous.
Proof.
(a) follows from ultratransitivity of action.
(b) $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})|_{G}=\mathfrak{E}_{\alpha_{G}}(\mathcal{U}_{X}^{G})$ and $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})|_{G}=\widetilde{\mathfrak{E}}_{\alpha_{G}}(\mathcal{U}_{X}^{G})$ in (b) follows from (a) and the definition of Ellis equiuniformity. $L\wedge R|_{G}=(L\wedge R)_{G}$ by Remark 3.16 and [37, Proposition 3.24] and Roelcke precompact. Since $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=L\wedge R$ (item (b) of Theorem 3.11), $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})|_{G}=(L\wedge R)_{G}$ and (b) is proved.
(c) Since $(L\wedge R)_{G}\subset\widetilde{\mathfrak{E}}_{\alpha_{G}}(\mathcal{U}_{X}^{G})\subset\widetilde{{\mathfrak{E}}}_{\star}((L\wedge R)_{G})$ by (b) and items (1), (2) of Theorem 3.2, $\widetilde{\mathfrak{E}}_{\alpha_{G}}(\mathcal{U}_{X}^{G})=\widetilde{{\mathfrak{E}}}_{\star}((L\wedge R)_{G})$ (item (f) of Theorem 2.22).
(d) follows from (b) and Remark 3.16.
$b_{r}G$ and $b_{r}\,\mbox{{\rm Aut}}(X)$ are uniformly isomorphic by the extension of the identity map on $G$ (since $G$ is dense in $\mbox{{\rm Aut}}\,(X)$).
If $X$ is continuous $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})=L\wedge R$ and $b_{r}\,\mbox{{\rm Aut}}(X)$ is a proper enveloping Ellis semigroup compactification (items (b’) and (c’) of Theorem 3.11).
If $b_{r}G$ is a proper enveloping Ellis semigroup compactification of $G$, then $b_{r}\,\mbox{{\rm Aut}}(X)$ is a proper enveloping Ellis semigroup compactification of $\mbox{{\rm Aut}}\,(X)$ ($b_{r}\,\mbox{{\rm Aut}}(X)$ is a right topological semigroup and the action $G\curvearrowright b_{r}\,\mbox{{\rm Aut}}(X)$ is extended to the action $\mbox{{\rm Aut}}(X)\curvearrowright b_{r}\,\mbox{{\rm Aut}}(X)$, because $G$ is dense in $\mbox{{\rm Aut}}(X)$ and $\mbox{{\rm Aut}}(X)$ is contained in the Raikov completion of $G$ to which the action is extended [31]). By Corollary 3.13 $X$ is continuous and the last statement is proved. ∎
Example 3.18.
The action of the Thompson group $(F,\tau_{\partial})$ is ultratransitive on the chain (not continuous) $\mathbb{Q}_{r}$ of dyadic rational numbers of (0,1)(0,1)( 0 , 1 ) and every automorphism from $F$ is a piecewise-linear map that have a finite number of breakpoints [10].
By Theorem 3.11 and Corollary 3.17 $F$ is Roelcke precompact, however, $L\wedge R\not\in\mathbb{E}(F)$ and the least proper enveloping Ellis semigroup compactification which is greater than the Roelcke compactification of $F$ is the completion of $(F,\widetilde{\mathfrak{E}}_{\alpha_{F}}(\mathcal{U}_{\mathbb{Q}_{r}}^{F}))$.
The description of $b_{m}\mathbb{Q}_{r}$ is given above. Clopen intervals with endpoints whose first coordinates (following the description) are dyadic rational numbers form the countable base of the topology of $b_{m}\mathbb{Q}_{r}$. Hence, $b_{m}\mathbb{Q}_{r}$ is a metrizable linearly ordered compacta (see, [33, Example 5.4]).
$b_{r}F=b_{r}\mbox{{\rm Aut}}\,(\mathbb{Q}_{r})$ by item (b) of Corollary 3.17 and its description is given in item (e) of Theorem 3.11.
3.2. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for the action of the unitary group on a Hilbert space
Let $U({\bf H})$ be the unitary group of a (complex or real) Hilbert space ${\bf H}$. It is a group of isometries for the action of $U({\bf H})$ on the unit sphere ${\rm S}\subset{\bf H}$, ${\rm S}$ is a coset space of $(U({\bf H}),\tau_{p})$ under its action [40, § 2.1] and the stabilizer $\mbox{{\rm St}}_{x}$ of any point $x\in{\rm S}$ is a neutral subgroup of $U({\bf H})$ [26, Remark 3.14], ${\rm B}$ in the unit ball in ${\bf H}$. The t.p.c. $\tau_{p}$ coincides with the strong operator topology (see for example, [39]) on $U({\bf H})$ (since for every $x\neq 0$ $||g(x)-x||<\varepsilon\Longleftrightarrow||g\big{(}\frac{x}{||x||}\big{)}-\frac{x}{||x||}||<\frac{\varepsilon}{||x||}$).
Let $\mathcal{L}_{\bf H}$ be the set of all finite-dimensional subspaces of ${\bf H}$, ${\rm B}_{L}=\{x\in L\ |\ ||x||\leq 1\}$ the unit ball, ${\rm S}_{L}=\{x\in L\ |\ ||x||=1\}$ is the unit sphere in $L$, $L\in\mathcal{L}_{\bf H}$.
Lemma 3.19.
The sets $V_{L,\varepsilon}=\{g\in U({\bf H})\ |\ ||g(x)-x||<\varepsilon,\ x\in{\rm B}_{L}\}$, $L\in\mathcal{L}_{{\bf H}}$, $\varepsilon>0$, form a base at the unit in $(U({\bf H}),\tau_{p})$.
Proof.
If $V=\{g\in U({\bf H})\ |\ ||g(x)-x||<\varepsilon,\ x\in\sigma\}\in N_{U({\bf H})}(e)$, where $\sigma=\{x_{1},\ldots,x_{n}\}\in\Sigma_{S}$ and $L={\rm span}(x_{1},\dots,x_{n})$ (the smallest linear subspace that contains $\sigma$), then $V_{L,\varepsilon}\subset V$.
For $V_{L,\varepsilon}$ let $e_{1},\ldots,e_{n}$ be the orthonormal basis of $L$. Then $V=\{g\in U({\bf H})\ |\ ||g(x)-x||<\varepsilon,\ x\in\sigma\}\subset V_{L,\varepsilon}$, where $\sigma=\{e_{1},\ldots,e_{n}\}$. From this lemma follows. ∎
Proposition 3.20.
The maximal equiuniformity $\mathcal{U}_{\rm S}$ on $S$ is totally bounded.
Proof.
Let $O\in N_{U({\bf H})}(e)$, $V\in N_{U({\bf H})}(e)$ is such that $V^{2}\subset O$, and, without loss of generality, is of the form $V=V_{L,\varepsilon}$, $L\in\mathcal{L}_{{\bf H}}$, $\varepsilon>0$, by Lemma 3.19, $\{Ox\ |\ x\in{\rm S}\}\in\mathcal{U}_{\rm S}$.
Let ${\bf H}={\rm L}\oplus{\rm L}^{\perp}$, $\mbox{{\rm pr}}:{\bf H}\to{\rm L}$ and $\mbox{{\rm pr}}_{\perp}:{\bf H}\to{\rm L}^{\perp}$ be the orthogonal projections, $\dim L=n$. One can define a continuous map $F:{\bf H}\to{\rm L}\times\mathbb{R}_{+}$, $F(x)=(\mbox{{\rm pr}}(x),||{\rm pr}_{\perp}(x)||)$. Since $||\mbox{{\rm pr}}(x)||^{2}+||\mbox{{\rm pr}}_{\perp}(x)||^{2}=1$ for $x\in{\rm S}$, the restriction $f=F|_{\rm S}$ of $F$ to the unit sphere sends points of ${\rm S}$ to the compact set ${\rm S}_{+}^{n}=\{(t,\tau)\in L\times\mathbb{R}_{+}:||t||^{2}+\tau^{2}=1\}$.
The map $f:{\rm S}\to{\rm S}_{+}^{n}$ is a composition of the projection $\mbox{{\rm pr}}|_{\rm S}:{\rm S}\to{\rm B}_{L}$ and the map $h:{\rm B}_{L}\to{\rm S}_{+}^{n}$, $h(x)=(x,\sqrt{1-||x||^{2}})$, which is a homeomorphism of ${\rm B}_{L}$ and ${\rm S}_{+}^{n}$. To show that the map $f$ is open it is sufficient to show that $\mbox{{\rm pr}}|_{\rm S}$ is open.
Let us note that the map $\sqrt{1-t^{2}}:[0,1]\to[0,1]$ is uniformly continuous. Hence, for any $b>0$ there exists $a>0$ such that if $|t-t^{\prime}|<a$, then $|\sqrt{1-t^{2}}-\sqrt{1-{t^{\prime}}^{2}}|<b$.
Take $x\in{\rm S}$ and its nbd $W(x)={\rm S}\cap W$, where $W=\{y\ |\ ||\mbox{{\rm pr}}(x)-{\rm pr}(y)||<a,||\mbox{{\rm pr}}_{\perp}(x)-\mbox{{\rm pr}}_{\perp}(y)||<b\}$. For any $x^{\prime}\in{\rm B}_{L}$ such that $||\mbox{{\rm pr}}(x)-x^{\prime}||<a$ take $y^{\prime}=\frac{\sqrt{1-||x^{\prime}||^{2}}}{||\mbox{{\rm pr}}_{\perp}(x)||}\mbox{{\rm pr}}_{\perp}(x)$. The point $(x^{\prime},y^{\prime})\in{\rm S}$ and
$||y^{\prime}-\mbox{{\rm pr}}_{\perp}(x)||=|\sqrt{1-||x^{\prime}||^{2}}-||\mbox{{\rm pr}}_{\perp}(x)|||=|\sqrt{1-||x^{\prime}||^{2}}-\sqrt{1-||\mbox{{\rm pr}}(x)||^{2}}|<b,$
since $|||\mbox{{\rm pr}}(x)||-||x^{\prime}|||\leq||\mbox{{\rm pr}}(x)-x^{\prime}||<a$ and the choice of $a$ and $b$. Thus, $\mbox{{\rm pr}}(W)\cap{\rm B}_{L}=\mbox{{\rm pr}}|_{\rm S}(W(x))$. This shows that $\mbox{{\rm pr}}|_{\rm S}$ (and, hence, $f$) is open.
For the action $\mbox{{\rm St}}_{L}\curvearrowright{\bf H}(={\rm L}\oplus{\rm L}^{\perp})$ the group $\mbox{{\rm St}}_{L}$ (stabilizer of points of the subspace $L$) is naturally identified with the unitary group $U({\rm L}^{\perp})$. The sphere ${\rm S}$ is an invariant subset which orbit space ${\rm S}/\mbox{{\rm St}}_{L}$ is identified with ${\rm S}_{+}^{n}$. Since the map of ${\rm S}$ onto the orbit space ${\rm S}/\mbox{{\rm St}}_{L}$ is open, ${\rm S}/\mbox{{\rm St}}_{L}$ is homeomorphic to ${\rm S}_{+}^{n}$.
The cover $\{Vx\ |\ x\in{\rm S}\}\in\mathcal{U}_{\rm S}$, its image $\{f(Vx)\ |\ x\in{\rm S}\}$ is an open cover of the compactum ${\rm S}/\mbox{{\rm St}}_{L}$ and has a finite refinement $f(V_{k})$, $k=1,\ldots,m$. Since $\mbox{{\rm St}}_{L}\subset V$, the finite open cover $f^{-1}(f(V_{k}))=\mbox{{\rm St}}_{L}V_{k}$, $k=1,\ldots,m$, of ${\rm S}$ is refined in $\{Ox\ |\ x\in{\rm S}\}$. Therefore, $\mathcal{U}_{\rm S}$ is totally bounded. ∎
Remark 3.21.
Another reasoning (which uses Roelcke precompactness of $U({\bf H})$ [45]) to prove that the maximal equiuniformity $\mathcal{U}_{\rm S}$ on $S$ is totally bounded can be found in [26, Example 3.31].
Lemma 3.22.
The base of the uniformity $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})$ is formed by the entourages
$\{(f,g)\in U({\bf H})^{2}\ |\ |(f(x),y)-(g(x),y)|<\epsilon,\ x,y\in{\rm B}_{L}\},\ L\in\mathcal{L}_{\bf H},$
where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the scalar product on ${\bf H}$.
Proof.
In [43, Corollary 2.2] it is shown that the unit ball ${\rm B}$ in ${\bf H}$ in the weak topology (the topology initial with respect to the projections onto finite-dimensional subspaces) is the maximal $G$-compactification of ${\rm S}$ (completion of $({\rm S},\mathcal{U}_{\rm S})$).
Therefore, the entourages from $\widetilde{\mathcal{U}}_{\rm S}$ are of the form
$\{(x,y)\in{\rm B}^{2}\ |\ ||\mbox{{\rm pr}}_{L}x-\mbox{{\rm pr}}_{L}y||<\epsilon\},\ L\in\mathcal{L}_{\bf H},$
and the entourages from $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})$ are of the form
$\{(f,g)\in U({\bf H})^{2}\ |\ ||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||<\epsilon,\ x\in\sigma\},\ \sigma\in\Sigma_{\rm B},\ L\in\mathcal{L}_{\bf H},$
where $\mbox{{\rm pr}}_{L}$ is the orthogonal projection of ${\bf H}$ onto $L$. Without loss of generality, one can assume that $\sigma=\{x_{1},\ldots,x_{n}\}$, $0\not\in\sigma$ and ${\rm span}\ (x_{1},\ldots,x_{n})=L$.
$||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||=||\mbox{{\rm pr}}_{L}(f(x)-g(x))||=\sup\limits_{z\in{\rm S}_{L}}|(f(x)-g(x),z)|=\sup\limits_{z\in{\rm S}_{L}}|(f(x),z)-(g(x),z)|.$
For $y\in{\rm B}_{L}$, $y\neq 0$
$|(f(x),y)-(g(x),y)|=||y||\cdot|(f(x),\frac{y}{||y||})-(g(x),\frac{y}{||y||})|\leq|(f(x),\frac{y}{||y||})-(g(x),\frac{y}{||y||})|.$
Therefore, $||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||=\sup\limits_{y\in{\rm B}_{L}}|(f(x),y)-(g(x),y)|$ and
$\{(f,g)\in U({\bf H})^{2}\ |\ |(f(x),y)-(g(x),y)|<\epsilon,\ x,y\in{\rm B}_{L}\}\subset$
$\subset\{(f,g)\in U({\bf H})^{2}\ |\ ||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||<\epsilon,\ x\in\sigma\}.$
For $\{(f,g)\in U({\bf H})^{2}\ |\ |(f(x),y)-(g(x),y)|<\epsilon,\ x,y\in{\rm B}_{L}\},\ L\in\mathcal{L}_{\bf H}$, let $e_{1},\ldots,e_{n}$ be the orthonormal basis of $L$, $\sigma=\{e_{1},\ldots,e_{n}\}$. For any $x\in{\rm B}_{L}$, one has $x=\lambda_{1}e_{1}+\ldots+\lambda_{n}e_{n}$, where $|\lambda_{k}|\leq 1$, $k=1,\ldots,n$, and
$|(f(x),y)-(g(x),y)|\leq\sum\limits_{k=1}^{n}|\lambda_{k}||(f(e_{k}),y)-(g(e_{k}),y)|\leq\sum\limits_{k=1}^{n}|(f(e_{k}),y)-(g(e_{k}),y)|.$
Therefore, if $||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||<\frac{\epsilon}{n}$, $x\in\sigma$, then $|(f(x),y)-(g(x),y)|\leq\varepsilon$ and
$\{(f,g)\in U({\bf H})^{2}\ |\ ||\mbox{{\rm pr}}_{L}f(x)-\mbox{{\rm pr}}_{L}g(x)||<\frac{\epsilon}{n},\ x\in\sigma\}\subset$
$\{(f,g)\in U({\bf H})^{2}\ |\ |(f(x),y)-(g(x),y)|<\epsilon,\ x,y\in{\rm B}_{L}\}.$
Theorem 3.23.
For the action $\alpha:U({\bf H})\times{\rm S}\to{\rm S}$ of the unitary group $U({\bf H})$ of a Hilbert space ${\bf H}$ on the the unit sphere ${\rm S}$ and the maps $\mathfrak{E}_{\alpha}:\mathbb{BU}({\rm S})\to\mathbb{BU}(U({\bf H}))$, $\widetilde{\mathfrak{E}}_{\alpha}:\mathbb{BU}({\rm S})\to\mathbb{E}(U({\bf H}))$ one has:
  • (a)
    $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})=\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})=L\wedge R=R_{\Sigma_{\rm S}}=R_{\Sigma_{\rm B}}$ (and, hence, $(U({\bf H}),\tau_{p})$ is Roelcke precompact [45], and $b_{r}U({\bf H})$ is a proper enveloping Ellis semigroup compactification),
  • (b)
    $b_{r}U({\bf H})$ is the space of all linear operators on ${\bf H}$ of norm 1$\leq$ 1≤ 1 in the weak operator topology [45],
  • (c)
    $b_{r}U({\bf H})$ is a semitopological semigroup [45].
Proof.
By Proposition 3.20 and item (1) of Theorem 3.2 to prove (a) one can check the fulfillment of the inclusions $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})\subset\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})$, $\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})\supset R_{\Sigma_{\rm S}}$ and $R_{\Sigma_{\rm S}}\supset R_{\Sigma_{\rm B}}$.
The proof of the inclusions $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})\subset\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})$ and $R_{\Sigma_{\rm S}}\supset R_{\Sigma_{\rm B}}$. For any $y\in{\rm B}$, $y\neq 0$, let $x=\frac{y}{||y||}\in{\rm S}$.
Then $\mbox{{\rm St}}_{y}=\mbox{{\rm St}}_{x}$ ($g(y)=y\Longleftrightarrow g(x)=x$, $g\in U({\bf H})$). Since $\mbox{{\rm St}}_{0}=U({\bf H})$, $R_{\Sigma_{\rm S}}=R_{\Sigma_{\rm B}}$.
If $f(x)\in Og(x)$, $O\in N_{U({\bf H})}(e)$, then there exists $h\in O$ such that $f(x)=(h\circ g)(x)$.
$f(y)=||y||\cdot f(x)=||y||\cdot(h\circ g)(x)=h(||y||\cdot g(x))=(h\circ g)(y).$
Therefore, if $f(x)$, $g(x)$ are in one element of the cover $\{Ot\ |\ t\in{\rm S}\}\in\mathcal{U}_{\rm S}$, then $f(y)$, $g(y)$ are in one element of its extension to ${\rm B}$. Since $f(0)=0$ for any $f\in U({\bf H})$, $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})=\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})$ (subbases of correspondent uniformities coincide).
The proof of the inclusion $\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})\supset R_{\Sigma_{\rm S}}$. In the proof of Theorem 1.1 from [45] (item (c)) it is shown that for any $V_{L,\varepsilon}$, $L\in\mathcal{L}_{{\bf H}}$, $\varepsilon>0$, there exist $\delta>0$, $v\in V_{L,\varepsilon}$ and $u\in\mbox{{\rm St}}_{L}$ such that if $f,g\in U({\bf H})$ and
[…]
then $f=ugv$.
The condition (w.o.t) defines the basic entourage of the uniformity, completion with respect to which is the (compact) unit ball in the space of all bounded operators equipped with the weak operator topology. This uniformity coincides with the uniformity $\widetilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{\rm S})=\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})$ by Lemma 3.22.
Let us note that if $f,g\in U({\bf H})$ and the condition (w.o.t) is fulfilled, then for the adjoint operators $f^{*},g^{*}\in U({\bf H})$ the same condition (w.o.t) is fulfilled. Indeed,
$|(f^{*}(x),y)-(g^{*}(x),y)|=|(x,f(y))-(x,f(y))|=|(f(y),x)-(g(y),x)|,\ \mbox{and}$
$|(f(y),x)-(g(y),x)|<\delta\Longleftrightarrow\ |(f(x),y)-(g(x),y)|<\delta,\ \forall\ x,y\in L,\ ||x||\leq 1,\ ||y||\leq 1.$
Therefore, if $f,g\in U({\bf H})$ and the condition (w.o.t) is fulfilled, then $f^{-1}=f^{*}=u^{\prime}g^{*}v^{\prime}$, where $u\in\mbox{{\rm St}}_{L}$, $v\in V_{L,\varepsilon}$. Hence, $f=v^{-1}gu^{-1}$, $u^{-1}\in\mbox{{\rm St}}_{L}$, $v^{-1}\in V_{L,\varepsilon}^{-1}=V_{L,\varepsilon}$ (since $||g^{-1}(x)-x||=||g(x)-x||$, $g\in U({\bf H})$) and, finally, $f\in V_{L,\varepsilon}g\mbox{{\rm St}}_{L}$. Hence, $\mathfrak{E}_{\alpha}(\mathcal{U}_{\rm S})\supset R_{\Sigma_{\rm S}}$.
[…]
For $f\in\mbox{{\rm cl}}~\tilde{\imath}(U({\bf H}))$ let us show that $f$ is the restriction of a linear map on ${\bf H}$.
(1) If $||x||,||y||\leq\frac{1}{2}$, then $f(x+y)=f(x)+f(y)$. Indeed, $||x+y||\leq||x||+||y||\leq 1$ and since $||f(x)||\leq||x||$, $x\in{\rm B}$, one has $||f(x+y)||\leq 1$, $||f(x)+f(y)||\leq 1$. Assume that $f(x+y)\neq f(x)+f(y)$. Take disjoint nbds $U$ of $f(x+y)$ and $V$ of $f(x)+f(y)$. Since addition in ${\bf H}$ is continuous there are nbds $Of(x)$ and $Of(y)$ such that $Of(x)+Of(y)\subset V$. Take the open set $O=[x,Of(x)]\cap[y,Of(y)]\cap[x+y,U]$ in ${\rm B}^{\rm B}$. $f\in O$, but $\tilde{\imath}(U({\bf H}))\cap O=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~\tilde{\imath}(U({\bf H}))$.
(2) If $||x||\leq\frac{1}{2}$, then $f(cx)=cf(x)$ for $c$ such that $|c|\leq 2$. Indeed, assume that $f(cx)\neq cf(x)$, $||x||\leq\frac{1}{2}$, $|c|\leq 2$, then $||cx||\leq 1$, $||cf(x)||\leq 1$. Take disjoint nbds $U$ of $f(cx)$ and $V$ of $cf(x)$. Since multiplication on the scalar in ${\bf H}$ is continuous there is nbd $Of(x)$ such that $cOf(x)\subset V$. Take the open set $O=[x,Of(x)]\cap[cx,U]$ in ${\rm B}^{\rm B}$. $f\in O$, but $\tilde{\imath}(U({\bf H}))\cap O=\emptyset$. Hence, $f\not\in\mbox{{\rm cl}}~\tilde{\imath}(U({\bf H}))$.
Put $\tilde{f}(x)=2||x||f(\frac{x}{2||x||})$, $x\in{\bf H}$. For any $c$ and any $x\in{\bf H}$
[…]
For any $x,y\in{\bf H}$
$\tilde{f}(x+y)\stackrel{{\scriptstyle(C)}}{{=}}2(||x||+||y||)f\big{(}\frac{x+y}{2(||x||+||y||)}\big{)}=2(||x||+||y||)\big{(}f\big{(}\frac{x}{2(||x||+||y||)}\big{)}+f\big{(}\frac{y}{2(||x||+||y||)}\big{)}\big{)}=$
$=2(||x||+||y||)f\big{(}\frac{x}{2(||x||+||y||)}\big{)}+2(||x||+||y||)f\big{(}\frac{y}{2(||x||+||y||)}\big{)}\stackrel{{\scriptstyle(C)}}{{=}}\tilde{f}(x)+\tilde{f}(y).$
From (2) and (C) $\tilde{f}(x)=f(x)$, $x\in{\rm B}$. Therefore any $f\in\mbox{{\rm cl}}~\tilde{\imath}(U({\bf H}))$ is a restriction of a linear map on ${\bf H}$.
Let $\tilde{f}$ be any linear operator on ${\bf H}$ of norm 1$\leq$ 1≤ 1, $f$ is its restriction to ${\rm B}$. The arbitrary nbd of $f$ is of the form $V=[x_{1},Of(x_{1})]\cap\ldots\cap[x_{n},Of(x_{n})]$, where, without loss of generality, $x_{1},\ldots,x_{n}\in{\bf H}$ are linearly independent, $Ox_{1},\ldots,Ox_{n}$ are open subsets of ${\rm B}$ in the weak topology such that ${\bf H}=L\oplus{\bf H}_{1}$ and $\mbox{{\rm pr}}^{-1}(\mbox{{\rm pr}}\ Of(x_{k}))=Of(x_{k})$, $k=1,\ldots,n$, where pr is the orthogonal projection of ${\bf H}$ to the finite dimensional subspace $L$.
Let $\varphi$ be any linear isometry of the ${\rm span}(x_{1},\dots,x_{n})$ to the $n$-dimensional subspace of ${\bf H_{1}}$ and $\tilde{\varphi}$ its extension to the linear isometry of ${\bf H}$ (any linear isometry of the orthogonal complement to the ${\rm span}(x_{1},\dots,x_{n})$ to the orthogonal complement to the image $\varphi({\rm span}(x_{1},\dots,x_{n}))$ can be chosen). Then $\tilde{\varphi}\in V$ and, hence, $f\in b_{r}U({\bf H})$.
(c) follows from (b) (all elements of $b_{r}U({\bf H})$ are continuous in the weak topology on ${\bf H}$). ∎